### v1

# Synthetic Series-Symbol Data Generation for Time Series Foundation Models

Abstract
Foundation models for time series analysis (TSA) have attracted significant attention. However, challenges such as training data scarcity and imbalance continue to hinder their development. Inspired by complex dynamic system theories, we design a series-symbol data generation mechanism, enabling the unrestricted creation of high-quality time series data paired with corresponding symbolic expressions. To leverage series-symbol data pairs with strong correlations, we develop SymTime, a pre-trained foundation model for enhancing time series representation using symbolic information. SymTime demonstrates competitive performance across five major TSA tasks when fine-tunes with downstream tasks, rivaling foundation models pre-trained on real-world datasets. This approach underscores the potential of series-symbol data generation and pretraining mechanisms in overcoming data scarcity and enhancing task performance. The code is available at https://github.com/wwhenxuan/SymTime.
1 Introduction
In recent years, with the rapid advancement of deep learning, foundation models for time series analysis (TSA) have garnered widespread attention due to their superior generalization capabilities, scalability and advantages in few-shot learning TimeMixer++ ; GPT4TS . Coupled with issues of data privacy DataPrivacy ; COMET , existing time series datasets are smaller compared to those in the fields of computer vision (CV) and natural language processing (NLP). Besides, current large-scale time series datasets face significant data imbalance issues, with certain types such as finance and healthcare still being relatively scarce (see Appendix B.4). According to scaling laws Neural-Sclaing-Laws , this can lead to performance bias in the time series foundation models, reducing their generalization capabilities on out-of-distribution data Time-Scaling-Laws ; Time-MoE .
To mitigate the issue of training data scarcity and imbalance, this paper, starting from Takens’ theorem Takens ; embedding , posits that time series are representations of complex dynamical systems TimeSeries4ComplexSystems ; li2023discover ; SNIP . Based on symbolic dynamics HAO1991161 , complex systems can be expressed abstractly using mathematical symbols and formulas Symbolic , with ordinary differential equations (ODE) and partial differential equations (PDE) being the most common methods for modeling complex systems DE4ModelingCS ; Differentiable . In an ideal scenario, continuously constructing diverse symbolic expressions allows us to cover a broader range of complex dynamical systems. As a result, the time series generated from these symbolic expressions exhibit rich and varied properties. To this end, we provide a series-symbol ($S^{2}$) dual-modality data generation mechanism. Simulation experiments demonstrate that this approach effectively mitigates the problem of training data scarcity. To encapsulate our work, the contributions are as follows:
  • Addressing Data Scarcity: Our approach overcomes the challenge of limited training data when building foundation models. The observation that the size of the $S^{2}$ dataset directly correlates with model performance on downstream tasks validates this point.
  • Introducing SymTime: We present SymTime, a scalable and efficient foundation model for time series analysis that leverages symbolic information to enhance representations. Pretrained on the constructed $S^{2}$ dataset, SymTime offers broader task generality compared to existing foundation models that are typically limited to zero-shot forecasting.
2 Related Work
Pre-trained foundation models (PTFMs) Time-LLM ; CLIP ; BERT ; BEiT have been demonstrated to adapt to a variety of downstream tasks after fine-tuning on specific datasets, exhibiting excellent generalization and scalability Chronos ; TimeMIL . Inspired by this, recent years have seen significant progress in PTFMs for TSA Transformer-in-TSA ; KDD-Survey , with the emergence of various pre-training methods. Moirai, through masked time series modeling (MTM) and reconstruction HiMTM ; SimMTM , has been pre-trained on large datasets ($27B$), yielding a universal forecasting model with zero-shot advantages MOIRAI . Timer, after generative pre-training on large datasets ($1B$), has performed well in forecasting Timer . TimeGPT trained a encoder-decoder Transformer with $100B$ data TimeGPT . COMET, using multi-level contrastive learning on a large ECG dataset, has obtained a medical time series PTFMs with few-shot advantages COMET .
As discussed in Appendix B.4, these baseline models still face challenges related to data scarcity and data imbalance. In the next section, we introduce the proposed data generation mechanism and the corresponding dual-modality foundation model designed to mitigate these issues. The review of other topics can be found in Appendix D.
3 Main Methods
Definition 1 Time Series Foundation Model.
It is a deep neural network pre-trained in a self‑supervised or unsupervised manner on large‑scale, diverse time series data. By learning generalizable time series representations, it can then rapidly adapt—via few‑shot or transfer learning—to efficiently solve a wide range of downstream time series tasks.
Theorem 1 Takens’ Theorem.
This theorem demonstrates that through phase space reconstruction Determining ; Characterization ; Differentiable ; Geometry , a univariate time series, as a low-dimensional projection of a high-dimensional complex system, can completely preserve the dynamic topology of the original system, thereby forming an effective external representation of the complex system Takens ; embedding ; Independent ; Nearest ; Predicting_multiple_observations ; spatiotemporal-dynamics ; Shalizi2006 ; probabilistic_takens .
Figure 1: The connection between time series and symbolic expressions (taking the Lorentz system as an example) kuznetsov2020lorenz .
Theorem 2 Symbolic Dynamics.
This theory encodes the evolution of continuous systems into finite symbolic expressions by discretizing the state space of complex systems, establishing an isomorphic relationship between symbolic expressions and system behaviors HAO1991161 ; sugihara1990 ; boaretto2021 ; tan2023 ; Lorentz . This means that any complex system can be modeled using symbolic expressions GA4SR1 ; GA4SR2 .
As show in Figure 1, the two aforementioned theorems, using complex dynamical systems as a conceptual bridge, fundamentally connect time series and symbolic expressions. They provide rigorous theoretical support for the semantic correspondence between temporal patterns and symbolic representations. Compared with previous time series data generation methods ForecastPFN ; Chronos , the $S^{2}$ data generation mechanism proposed in this paper is more in line with the nature of time series generation.
Figure 2: $S^{2}$ dataset generation mechanism (left) and SymTime network architecture (right).
3.1 Series-Symbol ($S^{2}$) Dataset Generation
The pre-training of SymTime relies on a large synthetic series-symbol ($S^{2}$) dataset [^1] The code for $S^{2}$ data generation is available at https://github.com/wwhenxuan/S2Generator.. The specific generation process is shown in Figure 2 (left). Firstly, we construct a multivariate input-output symbolic expression $f(\cdot)$ through random sampling DL4Symbolic . Then, we use the randomly generated sampling series $X\in\mathbb{R}^{M\times L}$ to forward propagate through the symbolic expression to obtain the generated series $Y=f(X)\in\mathbb{R}^{N\times L}$, where $N$ and $M$ represent the dimensions of the input and output series respectively, and $L$ is the length of the series. The mathematical symbols and their explanations in this section are shown in Table 9. In Appendix B.3, we present an analysis of the statistical characterization of the $S^{2}$ data. In Appendix B.7, we demonstrate that the time complexity of generating the $S^{2}$ data scales linearly with the series length $L$, approximating $\mathcal{O}(L)$.
3.1.1 Sampling of Functions
Figure 3: The process of building a binary tree when sampling symbolic expressions. (a) tree construction; (b) variable assignment to leaf nodes; (c) unary operator insertion.
Mathematical expressions can usually be represented using a tree structure, where constants and variables are the root nodes, binary operators are nodes with two children, and unary operators are nodes with one child. Therefore, we (1) build a binary tree over input variables using binary operators, (2) randomly insert constants and variables into the tree, and (3) add unary operator and perform affine transformation. The three key steps are illustrated in Figure 3.
Input/Output Dimension Selection.
Instead of randomly sampling input/output dimensions $M\sim\mathcal{U}(1,M_{\max})$ and $N\sim\mathcal{U}(1,N_{\max})$ as in prior work SNIP ; Symbolic ; neurosymbolic , we exhaustively traverse $M\in[1,M_{\max}],N\in[1,N_{\max}]$ (with $M_{\max}=6$, $N_{\max}=12$) to fully cover multivariate time series representations. An input dimension $M$ defines $M$ variable nodes $(x_{1},\dots,x_{M})$, while the output dimension yields $N$ generated series $y_{i}=f_{i}(x_{1},\dots,x_{M}),i=1,\dots,N$ DL4Symbolic .
Binary Operator Selection.
We sample the number of binary operators $b\sim\mathcal{U}(b_{\min},b_{\max})$ to define the root nodes of the expression tree. Each node’s operator is then drawn uniformly from $\mathcal{U}\{+,\,-$$\times\},$ enhancing the diversity and complexity of the generated expressions Symbolic ; SNIP ; GA4SR1 ; GA4SR2 .
Tree Construction and Leaf Assignment.
We randomly combine the $b$ binary operators to construct a binary tree to form the basic framework of the mathematical expressions (Figure 3a). Then, we randomly select $m$ variables from the $M$ set of variables $x_{1},\dots,x_{M}]$ and insert them into the symbolic expression consisting of binary operators, where $m\sim\mathcal{U}(1,M)$ (Figure 3b). If the inserted expression does not form a full binary tree, a random constant node is added to make it full.
Unary Operator Insertion.
After inserting the leaf nodes to form a complete binary tree, we select the number of unary operators $u$ from $\mathcal{U}(u_{\mathrm{min}},u_{\mathrm{max}})$ and insert unary operators at random positions in the binary tree. The available unary operators include $\{\mathrm{inv}$, $\mathrm{abs}$, $\mathrm{pow2},\mathrm{pow3},\mathrm{sqrt},\mathrm{sin},\mathrm{cos,tan,arctan,log,exp}\}$. This process is shown in Figure 3c. In Appendix B.8 we discussed the choice of unary operators.
Affine Transformation.
To further diversify the symbolic expressions, we perform random affine transformations on each random variable $x_{d}$ and unary operator $u_{d}$ in the binary tree. Specifically, we replace $x_{d}$ and $u_{d}$ with $ax_{d}+b$ and $au_{d}+b$, respectively, where $a$ and $b$ are random constants SNIP ; Symbolic . For example, $x_{1}\to ax_{1}+b$ and $\tan(\cdot)\to a\tan(\cdot)+b$ with contants $a,b$ sampled randomly.
3.1.2 Generating Inputs and Outputs Series
After obtaining symbolic expressions $f_{i}$, we sample $X\in\mathbb{R}^{M\times L}$ from mixed distributions  SNIP ; Symbolic ; DL4Symbolic and random-parameter ARMA($p,q$) processes ARIMA ; ARMA , then compute $Y\in\mathbb{R}^{N\times L},y_{i}=f_{i}(X)$. The ARMA($p$, $q$) model consists of moving average (MA) and autoregressive (AR) processes ARIMA_old , which can be expressed as:
${{Y}_{t}}={{\phi}_{1}}{{Y}_{t-1}}+{{\phi}_{2}}{{Y}_{t-2}}+\cdots+{{\phi}_{p}}{{Y}_{t-p}}+{{e}_{t}}-{{\theta}_{1}}{{e}_{t-1}}-{{\theta}_{2}}{{e}_{t-2}}-\cdots-{{\theta}_{q}}{{e}_{t-q}},$ (1)
where $p$ and $q$ represent the orders of the AR and MA models, respectively, $\phi_{p}$ and $\theta_{q}$ are the parameters of the AR and MA processes ARMA , and $e_{t}\sim\mathcal{N}(0,1)$ denotes the observed white noise sequence. Since ARMA possess both the temporal correlation of the AR process and the randomness of the MA process, series obtained from mixed distributions and ARMA sampling better reflect the characteristics of time series.
Sampling Strategy.
Each input series $X\in\mathbb{R}^{M\times L}$ is drawn either from a mixture of $k\sim\mathcal{U}(1,k_{\max})$ distributions (with weights $w_{j}\sim\mathcal{U}(0,1)$ normalized to $\sum_{j}w_{j}=1$, and each component chosen as $\mathcal{N}(\mu_{j},\sigma_{j}^{2})$, $\mu_{j}\sim\mathcal{N}(0,1)$, $\sigma_{j}\sim\mathcal{U}(0,1)$, or $\mathcal{U}(0,\mu_{j})$) with probability $P\leq 0.5$, or from an ARMA(p,q)(p,q) process ($p\sim\mathcal{U}(1,p_{\max})$, $q\sim\mathcal{U}(1,q_{\max})$, parameters $\phi_{i},\theta_{j}\sim\mathcal{U}(-1,1)$, and stationarity enforced by $\sum_{i}\phi_{i}<1,\;|\phi_{p}|<1$) otherwise.
Series Generation and Curation.
We normalize each $X$ per channel, compute $Y=f(X)$ via symbolic expressions, and discard any $X$ outside $f$’s domain or $|Y|>10^{4}$ SNIP ; Conformal . For each random seed, we traverse all input/output channels, sampling each expression once. The resulting $S^{2}$ dataset contains $40M$ series–symbol pairs ($50B$ total length); series are patched for the time series encoder PatchTST and expressions tokenized for the symbolic encoder DistilBERT .
3.2 Model Architecture of SymTime
As shown in Figure 2 (right), SymTime comprises a time series encoder, a symbol encoder, and momentum encoders, each trained with distinct objectives.
Time Series Encoder and Masked Time Series Modeling (MTM).
[…]
$\mathcal{L}_{\mathrm{mtm}}=\frac{1}{|\mathbf{M}_{T}|}\sum_{j\in\mathbf{M}_{T}}\|p_{j}-\hat{p}_{j}\|^{2},$ (2)
[…]
Symbolic Encoder and Masked Language Modeling (MLM).
We treat symbolic expression data as natural language and use the 6-layer DistilBERT DistilBERT as a symbol encoder to learn the representation of symbol through natural language mask modeling BERT . The loss optimized in this part is
$\mathcal{L}_{\mathrm{mlm}}=\frac{1}{|\mathbf{M}_{S}|}\sum_{j\in\mathbf{M}_{S}}\mathcal{H}\bigl(y_{j},\,p_{j}^{\mathrm{mask}}\bigr),$ (3)
[…]
Series–Symbol Contrastive Learning.
To leverage series-symbol data pairs with strong correlations, we employ contrastive learning to enhance time series representation using symbolic information. Using momentum encoders MoCo_v1 , we project [CLS] embeddings via linear projections $g_{t},g_{s}$ and define $\mathrm{sim}(t,s)=g_{t}(t_{\mathrm{cls}})^{\top}g^{\prime}_{s}(s^{\prime}_{\mathrm{cls}})$, where $g^{\prime}_{s}(s^{\prime}_{\text{cls}})$ is the normalized symbol features generated by the momentum model. Similarly, $\mathrm{sim}(s,t)=g_{s}(s_{\text{cls}})^{\mathrm{T}}g_{t}^{\prime}(t_{\text{cls}}^{\prime})$. We compute:
$p^{t2s}(t)=\frac{\exp(\mathrm{sim}(t,s_{m})/\tau)}{\sum_{m}\exp(\mathrm{sim}(t,s_{m})/\tau)},p^{s2t}(s)=\frac{\exp(\mathrm{sim}(s,t_{m})/\tau)}{\sum_{m}\exp(\mathrm{sim}(s,t_{m})/\tau)},$ (4)
where $\tau$ is a learnable temperature parameter COMET ; MoCo_v1 . Let $y^{t2s}(t)$ and $y^{s2t}(s)$ represent the one-hot similarity, with positive pairs having a probability of 1 and negative pairs having 0 MoCo_v1 . We optimize
$\mathcal{L}_{\mathrm{tsc}}=\tfrac{1}{2}\,\mathbb{E}\bigl[\mathcal{H}(y^{t2s},p^{t2s})+\mathcal{H}(y^{s2t},p^{s2t})\bigr].$ (5)
3.3 Momentum Distillation for Masked Data Learning
Inspired by ALBEF ALBEF , we treat random masking as noise and use momentum distillation to align the output representation of our encoder with its momentum counterpart. Let the similarity functions generated by the momentum encoders be $\text{sim}^{\prime}(t,s)=g_{t}(t^{\prime}_{\text{cls}})^{\mathrm{T}}g_{s}(s^{\prime}_{\text{cls}})$ and $\text{sim}^{\prime}(s,t)=g_{s}(s^{\prime}_{\text{cls}})^{\mathrm{T}}g_{t}(t^{\prime}_{\text{cls}})$. We compute soft pseudo targets $q^{t2s}(t)$ and $q^{s2t}(s)$ by replacing sim with $\text{sim}^{\prime}$ in Equation 4. In addition to the contrastive loss $\mathcal{L}_{\mathrm{tsc}}$ (Equation 5), we compute pseudo‐targets $q^{t2s},q^{s2t}$ from momentum‐encoder similarities $\mathrm{sim}^{\prime}$ and optimize
$\mathcal{L}_{\mathrm{tsc}}^{\mathrm{mod}}=\tfrac{1}{2}\ \mathbb{E}\left[\mathbf{KL}\left(q^{t2s}(t)\|p^{t2s}(t)\right)+\mathbf{KL}\left(q^{s2t}(s)\|p^{s2t}(s)\right)\right].$ (6)
The total pre-training objective is
$\mathcal{L}=\mathcal{L}_{\mathrm{mtm}}+\mathcal{L}_{\mathrm{mlm}}+\alpha\mathcal{L}_{\mathrm{tsc}}+(1-\alpha)\mathcal{L}_{\mathrm{tsc}}^{\mathrm{mod}}.$ (7)
3.4 Fine‑tuning for Downstream Tasks
We use the pre-trained time series encoder as backbone. After instance normalization ReVIN , we apply the following two strategies:
  • Classification: patch the series, encode, and classify via a linear head PatchTST ; TimeXer .
  • Reconstruction (forecasting, imputation, anomaly detection) Peri-midFormer : decompose each series into trend and periodic components; regress trend directly, patch and encode the periodic part, then recombine for the final output.
4 Experiments
We explore multiple representation measures and conduct experimental verification on a variety of downstream task datasets to answer the following key questions:
  • RQ1: Can the unrestrictedly generated $S^{2}$ dataset comprehensively cover diverse representation types of time series data?
  • RQ2: Can SymTime pre-trained on the $S^{2}$ dataset achieve competitive results across five major TSA tasks (forecasting, classification, imputation and anomaly detection)?
  • RQ3: Can SymTime learn fundamental representations of time series data on the synthetic $S^{2}$ dataset to alleviate the data scarcity in TSA?
  • RQ4: Are the multiple pre-training objectives in SymTime effective, and can symbol expressions enhance TSA task performance?
  • RQ5: How to demonstrate that SymTime learns semantic information of symbols?
4.1 Statistical Characterization and Representation Coverage of $S^{2}$ Dataset (RQ1)
Target.
We quantify the range of representations that the $S^{2}$ dataset can cover through statistical metrics (including stationarity (ADF Test) ADF , forecastability forecastable , frequency domain (FFT mean), seasonality seasonality , trend (Mann-Kendall Test) MK-test and prmutation entropy permutation ) (See Appendix B.5 for full descriptions).
(a) $100K$ single input channel
(b) $200K$ single input channel
(c) $100K$ single and dual channel
Figure 4: Radviz visualization of $S^{2}$ and Monash datasets.
Setup.
We use Radviz radviz to visualize high‑dimensional statistical features of 256256‑length time series segments from our synthetic $S^{2}$ and the Monash datasets Monash . From Monash (covering weather, traffic, electricity, tourism, medicine, and energy) we sample $200K$ segments per domain. For $S^{2}$, we sample 100K single‑channel segments (Figure 4a), then $200K$ single‑channel segments (Figure 4b), and finally $100K$ mixed single‑ and dual‑channel segments (Figure 4c).
Results.
Radviz visualization confirms that $S^{2}$ closely matches the Monash dataset across key statistics (stationarity, predictability, frequency, complexity, seasonality, trend), validating its use for pretraining. Expanding from $100K$ to $200K$ samples further broadens $S^{2}$’s coverage—surpassing Monash in some regions. Moreover, combining single- and dual-input samples dramatically increases diversity, as multi-variable expressions $f(x_{1},\cdots,x_{n})$ generate richer dynamics. These findings demonstrate that our infinitely scalable $S^{2}$ dataset covers the entire time series representation space.
4.2 Validation of SymTime in Five Time Series Analysis Tasks (RQ2)
Setup.
We pre-trained SymTime on the $50B$-scale $S^{2}$ dataset using Equation 7 as the pre-training objective, with the model architecture detailed in Table 16. Then, we evaluate SymTime on five TSA tasks: long-term forecasting, short-term forecasting, classification, imputation and anomaly detection, using the TimesNet benchmark TimesNet . We use mean squared error (MSE) and mean absolute error (MAE) as the metrics for long-term forecasting and imputation tasks; overall weighted average (OWA) for short-term forecasting, which is unique metrics for M4 benchmark M4team2018dataset ; accuracy for classification; precision, recall and F1 score for anomaly detection. Detailed descriptions of datasets and metrics for each task are provided in Appendix C.1 and C.2, while the pre-training and fine-tuning configurations for SymTime across downstream tasks are outlined in Appendix C.3 and C.4.
Baselines.
We compare with various baselines including Transformer-based models: PatchTST PatchTST , iTransformer iTransformer , Autoformer Autoformer , ETSformer ETSformer , FEDformer FEDformer , Non-stationary Transformer Non-stationary-transformers , Crossformer Crossformer , Informer Informer , Anomaly Transformer Anomaly-Transformer , Peri-midFormer Peri-midFormer ; LLM-based models: GPT4TS GPT4TS , Time-LLM Time-LLM , $S^{2}$IP-LLM S2IP-LLM ; CNN-based models: TimesNet TimesNet , TSLANet TSLANet , Rocket Rocket , InceptionTime (InTime) InceptionTime and MICN MICN ; MLP-based models: DLinear DLinear , LightTS LightTS , TimeMixer TimeMixer , FITS FITS and FilterNet FilterNet . We alse compare with the pre-trained foundation models: Moirai MOIRAI , Timer Timer , UniTS UniTS and Moment Moment . Some models can be applied to all 5 TSA tasks, while others are suitable for only one or some specific tasks. For those pre-trained foundation models, we first load their pre-trained parameters and then fine-tune them in the same way.
Figure 5: Model performance comparison with the state-of-the-art models in terms of five tasks (left). Complexity analysis on long time series forecasting tasks (ETTh1 dataset, forecasting length is 720 with 96 look-back windows) (right). Note that since the original backbone of Time-LLM Time-LLM has too many parameters, we replaced it with GPT2 GPT-2 .
Figure 6: Validation of SymTime in 5 time series analysis tasks. We only show the average results of all datasets in this figure. See Appendix A for full results and analysis on different tasks. To ensure fair experimentation, we use the original model hyperparameters. For long-term prediction tasks, when a model is tested at multiple look-back windows, we select the best result.
Main Results.
Figure 5 (left) compares SymTime with models of the same type, while Figure 6 presents SymTime’s performance against additional models across different tasks. These results demonstrate that SymTime, pre-trained on the $S^{2}$ dataset, successfully learns fundamental representations of time series data and achieves competitive results when fine-tuned on downstream tasks. For each different task, the specific experimental settings and results are shown in: long-term forecasting (Appendix A.1), short-term forecasting (Appendix A.2), classification (Appendix A.3), imputation (Appendix A.4) and anomaly detection (Appendix A.5).
Complexity Analysis.
We analyze the complexity of the model on the long-term forecasting ETTh1 dataset, with results shown in Figure 5 (right). We consider the parameter count, the GPU memory required for forward and backward propagation when the batch size is 1, Using MSE as an evaluation metric, we find that SymTime achieves better performance with a smaller model parameter count and memory capacity than existing foundation models in forecasting task.
4.3 The Impact of Pre-training Dataset Size on SymTime Performance (RQ3)
Table 1: Fine-tuning results of on long-term forecasting tasks with different pre-training dataset sizes. See Appendix F.8 for full results. The look-back window length for all experiments is 96. Red: best, Blue: second best.
{footnotesize }
Datasets ETTm1 ETTm2 ETTh1 ETTh2 Weather Electircity Traffic Exchange Average
Metrics MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
0B 0.401 0.409 0.293 0.339 0.487 0.474 0.376 0.412 0.257 0.289 0.193 0.284 0.471 0.310 0.383 0.415 0.358 0.366
1B 0.376 0.398 0.292 0.331 0.461 0.459 0.403 0.419 0.257 0.282 0.199 0.285 0.473 0.303 0.370 0.410 0.354 0.361
10B 0.376 0.393 0.281 0.329 0.444 0.444 0.376 0.408 0.250 0.279 0.196 0.286 0.473 0.294 0.368 0.407 0.345 0.355
25B 0.378 0.393 0.278 0.325 0.434 0.438 0.371 0.405 0.253 0.282 0.195 0.288 0.467 0.299 0.357 0.401 0.342 0.354
50B 0.371 0.390 0.274 0.321 0.430 0.436 0.365 0.402 0.247 0.276 0.187 0.276 0.457 0.291 0.359 0.401 0.336 0.349
Table 2: Fine-tuning results on short-term forecasting and imputation with different pre-training dataset sizes. See Appendix F.9 and F.11 for full results. Red: best, Blue: second best.
{footnotesize }
tasks Short-term Time Series Forecasting Time Series Imputation
Datasets Yearly Quartly Monthly Others Avg ETTm1 ETTm2 ETTh1 ETTh2 ECL Weather
Metrics Overall Weighted Average (OWA) MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
0B 0.782 0.913 0.964 1.097 0.887 0.042 0.122 0.038 0.106 0.112 0.230 0.065 0.160 0.058 0.155 0.036 0.053
1B 0.784 0.911 0.893 1.082 0.861 0.039 0.119 0.031 0.097 0.113 0.217 0.066 0.160 0.057 0.152 0.033 0.050
10B 0.783 0.905 0.896 1.055 0.859 0.038 0.119 0.030 0.095 0.107 0.213 0.063 0.158 0.056 0.151 0.033 0.048
25B 0.788 0.909 0.877 1.061 0.856 0.037 0.118 0.028 0.093 0.104 0.207 0.059 0.154 0.055 0.152 0.030 0.043
50B 0.786 0.872 0.872 1.045 0.849 0.036 0.117 0.026 0.088 0.095 0.201 0.058 0.148 0.054 0.151 0.028 0.038
(a) Time Series Classification
(b) Time Series Anomaly Detection
Figure 7: Fine-tuning results of on classification and anomaly detection tasks with different pre-training dataset sizes. We use critical difference diagrams to measure the performance of classification in (a) TSC-survey , where the specific values represent the comprehensive ranking of the model on multiple datasets. See Appendix F.10 and F.12 for full results.
Setup.
Following the scaling laws of neural networks Time-Scaling-Laws ; scaling-in-llm , a large-scale and representationally comprehensive pre-training dataset is critical for building foundation models. In Section 3.1 and 4.1, we introduce and validate the unrestricted generation capability and comprehensive representational coverage of the $S^{2}$ dataset. Building on this, we proportionally divide the generated $S^{2}$ dataset into timestamp-based subsets {$0B$, $1B$, $10B$, $25B$, 50B50B} and pre-train SymTime under identical configurations. To evaluate the effectiveness of pre-training, we fine-tune the models on five major TSA tasks (with $0B$ denoting direct fine-tuning without pre-training). The experimental results are summarized in Tables 1, Table 2 and Figure 7.
Results.
The results reveal a clear trend: as the scale of the pre-trained $S^{2}$ dataset increases, SymTime shows progressively enhanced performance across diverse downstream tasks. Furthermore, pre-trained SymTime significantly outperforms its non-pre-trained counterpart. Given the unrestricted generation capability and comprehensive representational coverage of our generation methods, the $S^{2}$ dataset can theoretically scale to infinite size, progressively enhancing SymTime’s downstream task performance through pre-training. This demonstrates that the proposed $S^{2}$ mechanism enables models to learn the representations of time series while effectively mitigating performance degradation caused by data scarcity and imbalanced distributions.
4.4 Ablation Study on Different Pre-training Objectives (RQ4)
Setup.
We conduct ablation studies on SymTime’s pre-training objectives using the ETTh1 and ETTh2 long-term forecasting datasets Informer . First, we establish 8 different control groups based on whether pre-training is performed, freezing the model and various pre-training losses: (1) Freeze, (2) Real-Data, (3) w/o Pre-train, (4) w/o MTM, (5) w/o MLM, (6) w/o T2S, (7) w/o S2T, (8) w/o Symbol and (9) w/o Distill. Specific explanations for the above control groups are provided in Appendix C.5 (Real-Data means we use real time series data of the same scale to pre-train the time series encoder only through the MTM.). Then, We adopt MSE as the metric, load the pre-trained parameters from each ablation configuration, and conduct multiple fine-tuning trials with varied random seeds on the ETT dataset. The average results with error are shown in Figure 8.
Figure 8: Ablation study on long-term forecasting task.
Results.
As shown in Figure 8, pre-training with standard configurations through Equation 7 significantly enhances SymTime’s long-term forecasting performance compared to control groups, demonstrating the validity of its pre-training objectives and the effectiveness of $S^{2}$ data generation. It is worth noting that, whether on real time series data (Real-Data) or on synthetic $S^{2}$ data (w/o Symbol), removing the symbol encoder part with contrastive learning and relying solely on MTM loss for temporal representation learning will degrade performance of SymTime. This suggests that semantic information provided by the symbolic encoder and series-symbol contrastive learning improves the temporal encoder’s capability in long-term forecasting AutoTimes . Similar findings are replicated in short-term forecasting tasks, as detailed in Appendix C.6.
The Backbone in SymTime.
There is a time series encoder with a Transformer architecture and a symbol encoder consisting of a pre-trained LLM in SymTime. In Appendix C.7, we perform ablation experiments on the model architecture by replacing the backbone of SymTime. The results are shown in Figure 17. It is clear that replacing the backbone has no significant impact on downstream tasks. The improvement in model performance mainly comes from the dual-modal pre-training paradigm of time series and symbolic expressions (Equation 7).
4.5 Series-Symbol Representation Learning (RQ5)
We further analyze the data representations learned by SymTime through masked modeling and cross-modal contrastive loss. Given that contrastive learning aligns mutually positive samples in the representation space SimCLR ; MoCo_v1 , we pre-train SymTime on the $S^{2}$ dataset and extract its time series encoder and symbol encoder. Using generated univariate time series and unary symbolic expressions, we examine whether the representation spaces of encoders evolve during pre-training. Additionally, we specifically evaluate the time series encoder’s proficiency in temporal representation learning. Results demonstrate that the time series encoder, pre-trained via large-scale masked modeling PatchTST ; SimMTM , achieves zero-shot imputation capability on both the $S^{2}$ dataset and real-world time series data. Detailed experimental analyses are provided in Appendix B.6.
4.5.1 Time Series Encoder Representation
Setup.
We sample $20K$ single-channels series–symbol pairs from $S^{2}$, with each time series labeled by one unary operator (e.g., inv, sin, log). After patching, the time series segments are encoded by time series encoder of SymTime with and without series-symbol contrastive pretraining. The output embeddings are reduced via t-SNE t-SNE .
Results.
As shown in Figure 9 (a)(b), The untrained encoder produces entangled clusters (aside from outliers like inv, exp), whereas the pretrained encoder forms clear, operator-specific clusters (trigonometric: sin, cos; polynomial: pow2, pow3, sqrt; etc.), confirming that contrastive learning aligns symbolic semantics with series representations. The change in representation space before and after pre-training also proves that the time series encoder has mastered the semantic information of symbolic expressions. Ablations show both contrastive losses are essential.
4.5.2 Symbol Encoder Representation
Setup.
We also use $20K$ series-symbol pairs to verify the representation space change of the symbol encoder. First, we tokenize the symbolic expression. Then, we input it into the encoder to obtain its [CLS] token CLIP ; ALBEF and visualize it using the t-SNE t-SNE .
Results.
As shown in Figure 9 (c)(d), similar to the time series encoder, the pre-trained DistilBERT-based symbol encoder initially struggles to distinguish between different types of symbol DistilBERT . However, after pre-training with MLM and contrastive learning, the encoder forms distinct clusters in the representation space for symbolic expressions of the same type SNIP . Furthermore, the paired time-series data and their corresponding symbolic expression exhibit similar clustering characteristics, as shown in Figure 9 (b) and (d). This demonstrates that cross-modal representation learning effectively brings semantically related data points closer together in the representation space.
(a) without pre-training
(b) with pre-training
(c) without pre-training
(d) with pre-training
Figure 9: The t-SNE visualization of time series encoder and symbol encoder in SymTime representation space with 50 perplexity. (a)(b) time series encoder; (c)(d) symbol encoder.
5 Conclusion
To mitigate the challenges of data scarcity and distribution imbalance in time series analysis, we introduce a dual-modality series-symbol ($S^{2}$) data generation mechanism that enables the unrestricted creation of high-quality time series data, along with corresponding symbolic representations. Leveraging this large-scale series-symbol synthetic dataset, we propose SymTime, a pre-trained foundation model that integrates both time series representations and symbolic semantic information through masked modeling and contrastive learning. Our pre-trained model demonstrates exceptional performance across five major TSA tasks, highlighting the effectiveness of both our data generation strategy and pre-training methodology. Looking ahead, we aim to scale up our approach by training larger models on synthetic datasets, further boosting performance on downstream tasks.
Acknowledgement
This work was supported in part by Natural Science Basic Research Program of Shaanxi under Grant 2025JC-QYCX-060, in part by the National Natural Science Foundation of China under Grant 62206205, 62201419 and 62471371, in part by the Young Talent Fund of Association for Science and Technology in Shaanxi, China under Grant 20230129, in part by the Guangdong High-level Innovation Research Institution Project under Grant 2021B0909050008, in part by the Guangzhou Key Research and Development Program under Grant 202206030003, and in part by the Fundamental Research Funds for the Central Universities under Grant QTZX25111. This research was supported by Huawei’s Al Hundred Schools Program and was carried out using the Huawei Ascend AI technology stack.
Appendix A Main Results and Conclusions of the Five Tasks in Time Series Analysis
A.1 Long-term Forecasting
Setup.
Time series forecasting, which analyzes historical data patterns to predict future trends, is crucial for financial market analysis, inventory management, energy demand and other fields TimeMixer++ ; Timer-XL . We adopt 8 real-world benchmark datasets for long-term forecasting, including ETTm1, ETTm2, ETTh1, ETTh2 Informer , Weather weather , ECL ECL , Traffic traffic and Exchange LSTNet . The forecasting lengths are set to $\{96,192,336,720\}$. To ensure the fairness of the experiment, we set up three different look-back windows {$96$, $336$, 512512} for the experiment, except Moirai and Timer are $672$. For different models, we try our best to ensure that they are tested according to their original experimental configuration. For foundation models that can perform zero-shot forecasting and have been pre-trained, such as Moirai MOIRAI , Timer Timer , Moment Moment and Chronos Chronos , we first load the pre-trained parameters of the model, and then perform supervised fine-tuning on it in the same way.
Table 3: Long-term forecasting task with 96 look-back windows. The results are averaged from four different series length {96,192,336,72096,192,336,720}. (* means former.) See Appendix F.1 for full results. Red: best, Blue: second best. The standard deviation is within 1%.
{footnotesize }
SymTime Peri-mid* Moriai Timer Time-LLM TSLANet $S^{2}$IP-LLM GPT4TS TimeMixer
Model (Ours) Peri-midFormer MOIRAI Timer Time-LLM TSLANet S2IP-LLM GPT4TS TimeMixer
Metrics MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTm1 0.371 0.387 0.409 0.410 0.398 0.417 0.388 0.402 0.369 0.394 0.377 0.397 0.374 0.404 0.369 0.395 0.382 0.397
ETTm2 0.274 0.321 0.290 0.328 0.296 0.348 0.405 0.408 0.275 0.324 0.283 0.327 0.266 0.325 0.264 0.328 0.279 0.325
ETTh1 0.430 0.436 0.455 0.446 0.441 0.454 0.434 0.444 0.438 0.445 0.448 0.441 0.456 0.454 0.434 0.440 0.453 0.441
ETTh2 0.365 0.402 0.400 0.416 0.402 0.411 0.428 0.441 0.369 0.407 0.355 0.391 0.362 0.405 0.359 0.403 0.388 0.408
Weather 0.247 0.276 0.262 0.283 0.265 0.299 0.329 0.358 0.247 0.269 0.259 0.352 0.243 0.274 0.265 0.285 0.253 0.280
ECL 0.187 0.276 0.178 0.267 0.167 0.252 0.177 0.267 0.180 0.269 0.199 0.283 0.191 0.283 0.206 0.291 0.185 0.274
Traffic 0.457 0.291 0.458 0.295 0.424 0.289 0.436 0.284 0.471 0.334 0.463 0.310 0.417 0.306 0.491 0.320 0.499 0.306
Exchange 0.359 0.401 0.388 0.417 0.373 0.417 0.382 0.425 0.376 0.414 0.368 0.414 0.472 0.478 0.370 0.411 0.403 0.423
Average 0.336 0.349 0.355 0.358 0.346 0.361 0.372 0.378 0.341 0.357 0.344 0.364 0.348 0.366 0.345 0.359 0.355 0.357
Table 4: Long-term forecasting task with 336 look-back windows. The results are averaged from four different series length {96,192,336,72096,192,336,720}. See Appendix F.2 for full results. Red: best, Blue: second best. The standard deviation is within 1%.
{footnotesize }
SymTime PatchTST TimeMixer TimesNet Autoformer DLinear iTransformer TimeXer FEDformer
Model (Ours) PatchTST TimeMixer TimesNet Autoformer DLinear iTransformer TimeXer FEDformer
Metrics MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTm1 0.350 0.382 0.352 0.382 0.368 0.392 0.421 0.423 0.618 0.539 0.357 0.379 0.368 0.395 0.372 0.395 0.441 0.452
ETTm2 0.256 0.316 0.258 0.315 0.262 0.318 0.282 0.334 0.400 0.420 0.291 0.353 0.272 0.329 0.262 0.317 0.325 0.377
ETTh1 0.413 0.432 0.419 0.432 0.430 0.437 0.485 0.480 0.580 0.539 0.425 0.440 0.450 0.457 0.493 0.483 0.450 0.472
ETTh2 0.341 0.390 0.331 0.379 0.396 0.425 0.409 0.440 0.663 0.604 0.490 0.476 0.390 0.416 0.375 0.410 0.430 0.467
Weather 0.238 0.273 0.258 0.292 0.235 0.273 0.250 0.286 0.441 0.450 0.245 0.298 0.238 0.272 0.287 0.290 0.313 0.363
ECL 0.164 0.258 0.165 0.294 0.169 0.260 0.197 0.297 0.236 0.346 0.170 0.269 0.163 0.257 0.172 0.267 0.214 0.327
Traffic 0.391 0.267 0.396 0.268 0.411 0.271 0.615 0.331 0.676 0.413 0.465 0.320 0.401 0.283 0.452 0.281 0.610 0.376
Exchange 0.367 0.406 0.385 0.420 0.415 0.438 0.548 0.532 1.053 0.807 0.448 0.462 0.392 0.427 0.409 0.500 0.376 0.427
Average 0.315 0.341 0.320 0.348 0.336 0.352 0.401 0.390 0.583 0.515 0.361 0.375 0.347 0.355 0.353 0.368 0.395 0.408
Table 5: Long-term forecasting task with 512 look-back windows. The results are averaged from four different series length {96,192,336,72096,192,336,720}. See Appendix F.3 for full results. Red: best, Blue: second best. The standard deviation is within 1%.
{footnotesize }
SymTime PatchTST TimeMixer TimesNet Autoformer DLinear iTransformer TimeXer FITS
Models Ours PatchTST TimeMixer TimesNet Autoformer DLinear iTransformer TimeXer FITS
Metrics MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTm1 0.356 0.380 0.352 0.382 0.371 0.392 0.425 0.430 0.556 0.518 0.358 0.380 0.367 0.397 0.378 0.401 0.374 0.384
ETTm2 0.265 0.320 0.256 0.317 0.263 0.323 0.294 0.344 0.371 0.416 0.275 0.340 0.273 0.331 0.274 0.329 0.254 0.313
ETTh1 0.414 0.432 0.413 0.434 0.429 0.444 0.481 0.486 0.627 0.579 0.418 0.438 0.446 0.460 0.475 0.479 0.418 0.441
ETTh2 0.365 0.405 0.357 0.409 0.373 0.410 0.397 0.432 0.687 0.609 0.499 0.478 0.388 0.417 0.354 0.400 0.363 0.408
weather 0.234 0.273 0.245 0.284 0.231 0.271 0.251 0.288 0.489 0.486 0.241 0.292 0.249 0.280 0.282 0.284 0.244 0.281
ECL 0.163 0.267 0.169 0.269 0.177 0.274 0.201 0.302 0.353 0.393 0.167 0.267 0.162 0.257 0.171 0.270 0.169 0.265
Traffic 0.395 0.268 0.399 0.272 0.410 0.270 0.624 0.334 0.705 0.435 0.433 0.305 0.383 0.273 0.466 0.287 0.420 0.287
Exchange 0.384 0.412 0.398 0.423 0.517 0.497 0.718 0.608 0.944 0.768 0.500 0.494 0.427 0.467 0.514 0.506 0.393 0.439
Average 0.322 0.345 0.324 0.349 0.346 0.360 0.424 0.403 0.591 0.525 0.361 0.374 0.337 0.360 0.364 0.369 0.329 0.352
Results.
Tables 3, Table 4, and Table 5 respectively show the results of SymTime with look-back windows of length {$96$, $336$, 512512}. It can be seen that SymTime demonstrates excellent performance in time series long-term forecasting tasks. Our model surpasses Peri-midFormer, GPT4TS and TimesNet, which are foundation models for the five major tasks, as well as Moirai and Timer, two general forecasting models. Compared with Time-LLM and $S^{2}$IP-LLM, which use pre-trained large language models as backbone, SymTime achieves better results with a more lightweight Transformer encoder.
A.2 Short-term Forecasting
Setup.
We adopt M4 benchmark M4team2018dataset for short-term forecasting, which contains the yearly, quarterly and monthly collected univariate marketing data. Then, we use symmetric mean absolute error (SMAPE), mean absolute scaled error (MASE) and overall weighted average (OWA) to measure the forecasting performance, which are calculated as detailed in Appendix C.2.
Results.
Table 6 indicates that SymTime after pre-training, surpasses TimeMixer, Peri-midFormer and TimesNet on the short-term forecasting tasks in terms of SMAPE, MASE and OWA metrics, achieving state-of-the-art performance. Specifically, SymTime performs well on Yearly, Quarterly and Monthly datasets, demonstrating its capability to capture not only the trends of annual variations but also the cyclic characteristics of seasonal and monthly encoding.
Table 6: Short-term forecasting task on M4. The prediction lenghs are $\{6,48\}$ and results are weighted averaged from several datasets under different sample intervals. (* means former, TMixer is TimeMixer, $S$-LLM is $S^{2}$IP-LLM, T-LLM is Time-LLM). See Appendix F.4 for full results. Red: best, Blue: second best. The standard deviation is within 1%.
{footnotesize }
SymTime Peri-mid* $S$-LLM T-LLM GPT4TS TMixer PatchTST iTrans* TimesNet DLinear LightTS FED* In*
Models (Ours) Peri-midFormer S2IP-LLM Time-LLM GPT4TS TimeMixer PatchTST iTransformer TimesNet DLinear LightTS FEDformer Informer
SMAPE 11.785 11.897 12.514 12.584 12.367 11.885 12.866 13.233 11.888 12.500 11.962 12.605 15.018
MASE 1.584 1.607 1.726 1.763 1.767 1.598 1.734 1.850 1.607 1.678 1.609 1.677 2.096
OWA 0.849 0.859 0.913 0.915 0.918 0.856 0.928 0.972 0.858 0.899 0.862 0.903 1.102
A.3 Classification
Setup.
Time series classification is crucial for the identification and diagnosis of patterns in complex systems and plays a significant role in various fields such as financial analysis, medical diagnosis and industrial monitoring TSC-survey . Using the experimental setup from TimesNet TimesNet , we test SymTime’s discriminative ability on 10 UEA multivariate time series classification datasets UEA , including categories such as Industry, Face Detection, ECG, Voice and Transportation.
Figure 10: Comparison of the average accuracy of SymTime and other baselines on 10 UEA datasets. See Appendix F.5 for full results.
Results.
As shown in Figure 10, SymTime achieves an average accuracy of 74.9%, surpassing all baselines, indicating that SymTime is competitive in classification tasks.
A.4 Imputation
Setup.
Sensors monitoring complex systems in the real world may experience distortions or malfunctions, leading to partial missing data in the collected time series. Therefore, time series imputation is crucial for the recovery of complete datasets. We verify SymTime’s imputation capabilities on 6 datasets: ETTm1, ETTm2, ETTh1, ETTh2 Informer , Weather weather and ECL ECL . To test the model’s imputation ability under varying degrees of missing data, we add random masks at proportions of $12.5\%,25\%,37.5\%,50\%$} in point level on time series of length 96. Since SymTime was pre-trained by randomly masking patches level for series reconstruction and masks are added randomly in point level in the imputation task. Considering the differences between these masking approaches and the potential disruption of the series’s original trends and periodic features at higher mask rates, we adopt per-interpolation for the masked series from Peri-midFormer . Analysis and ablation experiments regarding this method are presented in Appendix C.8. The results demonstrate that per-interpolation can be used as a model-independent feature engineering to improve the performance of downstream tasks.
Table 7: Imputation task, where we randomly mask {12.5%, 25%, 37.5%, 50%} time points of length-96 time series. The reuslts averaged from 4 different mask ratios. (* means former.) See Appendix F.6 for full results. Red: best, Blue: second best.
{footnotesize }
SymTime GPT4TS TimesNet Peri-mid* Moment iTrans* PatchTST DLinear LightTS
Model (Ours) GPT4TS TimesNet Peri-midFormer Moment iTransformer PatchTST DLinear LightTS
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTm1 0.036 0.116 0.028 0.109 0.027 0.107 0.036 0.116 0.074 0.168 0.072 0.185 0.049 0.143 0.090 0.204 0.068 0.182
ETTm2 0.026 0.088 0.022 0.088 0.022 0.089 0.026 0.087 0.031 0.108 0.082 0.191 0.030 0.101 0.102 0.212 0.068 0.176
ETTh1 0.095 0.201 0.093 0.200 0.089 0.199 0.091 0.196 0.139 0.234 0.148 0.269 0.126 0.231 0.169 0.283 0.159 0.278
ETTh2 0.058 0.148 0.052 0.147 0.050 0.148 0.057 0.147 0.061 0.159 0.139 0.254 0.066 0.164 0.163 0.273 0.143 0.258
ECL 0.054 0.151 0.093 0.212 0.094 0.211 0.063 0.169 0.094 0.211 0.099 0.224 0.078 0.192 0.128 0.256 0.108 0.238
Weather 0.028 0.038 0.032 0.058 0.030 0.056 0.029 0.041 0.035 0.075 0.052 0.114 0.033 0.057 0.053 0.116 0.047 0.106
Average 0.049 0.124 0.053 0.136 0.052 0.135 0.050 0.126 0.072 0.159 0.099 0.206 0.064 0.148 0.118 0.224 0.099 0.206
Results.
Table 7 shows that SymTime outperforms Peri-midFormer, GPT4TS and TimesNet in overall performance establishing SymTime as the latest state-of-the-art approach. Although SymTime’s performance on the ETT series of datasets is not as strong as GPT4TS, it achieves more significant effects on datasets with a higher number of channels, such as ECL and Weather.
A.5 Anomaly Detection
Table 8: Anomaly detection task, where we calculate the F1-score (as %) for each dataset. (* means former, TNet is TimesNet, PTST is PatchTST.) A higher value of F1-score indicates a better performance. See Appendix F.7 for full results. Red: best, Blue: second best.
{footnotesize }
SymTime UniTS Peri-mid* GPT4TS TNet PTST LightTS DLinear iTrans* Anomaly Stationary Cross* In* Auto*
Model (Ours) UniTS Peri-midFormer GPT4TS TimesNet PatchTST LightTS DLinear iTransformer Anomaly-Transformer Non-stationary-transformers Crossformer Informer Autoformer
SMD 85.66 83.69 84.08 84.49 84.37 84.62 82.53 79.76 80.19 85.49 82.97 77.22 77.88 71.17
MSL 81.77 81.16 80.68 82.03 81.14 78.70 78.95 81.87 72.47 83.31 76.68 80.59 81.07 82.22
SMAP 73.43 74.00 67.53 68.85 69.05 68.82 69.21 67.30 66.72 71.18 69.02 67.12 73.26 73.97
SWaT 93.61 92.51 91.64 92.60 92.61 85.72 93.33 92.66 92.64 83.10 92.24 90.22 80.35 79.19
PSM 97.10 97.31 96.21 97.09 97.06 96.08 97.15 96.64 94.88 79.40 97.23 92.52 90.43 88.24
Avg F1 86.31 85.73 84.03 85.01 84.85 82.79 84.23 83.64 81.38 80.50 83.63 81.53 80.60 78.96
Setup.
Time series anomaly detection is crucial for rapidly identifying anomalies in critical areas, aiding in risk prevention and decision optimization. Due to the difficulty in annotating time series anomalies, we focus primarily on unsupervised anomaly detection. We conduct experiments on 5 widely used anomaly detection datasets: SMD and SMAP SMD , MSL MSL , SWaT SWaT , PSM PSM , encompassing service monitoring, space & earth exploration, and water treatment applications. We adopt the same data preprocessing method as the Anomaly Transformer Anomaly-Transformer , dividing the data into non-overlapping segments of length 100 for reconstruction. Specifically, normal data is used for model training and we employ a simple reconstruction loss to help the model learn the distribution of normal data Peri-midFormer . In subsequent testing phases, reconstructed outputs exceeding a specified threshold are considered anomalies.
Results.
The results in Table 8 show that SymTime surpasses all previous models such as TimesNet and GPT4TS in the time series anomaly detection task and becomes the latest SOTA model. Compared with UniTS UniTS pre-trained on real time series data, SymTime is pre-trained on synthetic datasets and achieves better model performance.
Appendix B Analysis of Series-Symbol ($S^{2}$) Dataset and Model Pre-training
Table 9: Some symbols used in data generation and their explanations.
Symbols Explanation Symbols Explanation
$X$ sampling series $Y$ generated series
$f(\cdot)$ symbolic expression $e_{t}$ white noise sequence
$M$ the input channels number NN the output channels number
$\mathcal{U}$ uniform distribution $\mathcal{N}$ normal distribution
$P$ probability of selecting sampling methods $k$ total number of mixed distributions
$p$ the order of the AR process $q$ the order of the MA process
$\phi_{p}$ the parameters of AR process $\theta_{q}$ the parameter of MA process
Based on the viewpoint of complex dynamic system modeling, this paper proposes a bimodal series-symbol ($S^{2}$) generation mechanism to alleviate the problem of shortage of training data in the field of time series analysis. Table 9 shows the specific symbols we used in constructing symbolic expressions in Section 3.1. This part is mainly divided into the following sections:
  • B.1 Series-Symbol Data Display: This section shows the symbolic expressions and time series data that can be generated through the $S^{2}$ data generation mechanism.
  • B.2 Composition and Usage of the Series-Symbol Dataset: This section introduces the specific composition of the $S^{2}$ dataset and how to use it when training SymTime.
  • B.3 Statistics Analysis: This section examines and analyzes the basic statistical representation of the $S^{2}$ data.
  • B.4 Analysis of Existing Large-scale Datasets for Time Series Pre-training: This section analyzes the data shortage and distribution imbalance faced by large-scale time series pre-training datasets in TSA.
  • B.5 $S^{2}$ Dataset Statistical Characterization Coverage Experiments: This section details the specific configuration of the characterization coverage experiments in the Section 4.1.
  • B.6 Masked Time Series Modeling and Zero-shot Imputation for Representation Learning:
  • B.7 Time Complexity Analysis of $S^{2}$ Data Generation Mechanism: In this section we analyze in detail the time complexity of the $S^{2}$ data generation mechanism.
  • B.8 The Selection of the Unary Operators: In this section we analyze the unary operators used in $S^{2}$ data generation.
  • B.9 The Limitations of $S^{2}$ Generation Mechanism.
B.1 Series-Symbol Data Display
In Figure 11, we show the visualization of the generated series from 1 input channel and 1 output channel to 4 input channels and 4 output channels. We show two sets of cases for each input and output channel. The symbolic expressions $f(\cdot)$ for the generated series in (a), (c), (e) and (g) in Figure 11 are shown above.
The symbolic expressions with text format are shown as follow: Symbolic expression of Figure 11 (a) $y_{1}$ = (-0.795 add ((-0.675 mul ((0.999 add (-6.7 mul $x_{1}$)))**2) add ((-0.798 mul inv((-5.99 add (-0.751 mul $x_{1}$)))) sub (9.68 mul sqrt((-7.37 add (0.756 mul $x_{1}$))))))) Symbolic expression of Figure 11 (c) $y_{1}$ = (-3.39 add (((0.56 mul (inv((-98.9 add (58.2 mul $x_{2}$))) mul ((-19.7000 mul $x_{1}$) sub (31.9000 mul $x_{2}$)))) sub (40.4000 mul $x_{1}$)) add (0.71 mul (((7.13 mul $x_{2}$) sub (-1.68 mul ($x_{1}$ mul sqrt((-92.8000 add (0.054 mul ($x_{2}$ mul ((0.327 mul $x_{2}$) sub (2.3 mul $x_{2}$))))))))) mul $x_{1}$)))) $y_{2}$ = (1.0 add ((68.9 mul $x_{2}$) sub (((80.9 mul ($x_{1}$ mul ($x_{1}$ mul ((6.1000 mul $x_{2}$) sub ((34.2 mul sqrt((64.4 add (29.2000 mul $x_{1}$)))) add (-5.24 mul $x_{1}$)))))) add (6.78 mul $x_{2}$)) sub (((0.5730 mul $x_{1}$) sub ((2.34 mul $x_{2}$) sub (-6.72 mul $x_{1}$))) add (0.966 mul sqrt((76.8000 add (-7.79 mul $x_{1}$)))))))) Symbolic expression of Figure 11 (e) $y_{1}$ = (0.795 add ((0.42 mul $x_{3}$) sub ((4.39 mul $x_{1}$) add (((0.1430 mul $x_{2}$) sub ((-5.28 mul $x_{3}$) add (((-0.028 mul ((((1.27 mul $x_{3}$) sub (((0.331 mul $x_{2}$) sub ((2.99 mul $x_{3}$) add (-0.932 mul (((0.606 mul $x_{1}$) sub (0.967 mul $x_{3}$)) mul $x_{3}$)))) sub (-0.609 mul $x_{3}$))) add (-1.25 mul $x_{1}$)) mul $x_{1}$)) sub (77.3000 mul $x_{1}$)) sub (1.93 mul $x_{3}$)))) sub (16.7 mul $x_{3}$))))) $y_{2}$ = (-9.2900 add ((0.398 mul ((((-49.7 mul $x_{1}$) sub ((5.93 mul sin((6.54 add (-0.045 mul $x_{1}$)))) add ((62.3000 mul inv(((0.138 mul $x_{2}$) add (29.0 mul $x_{1}$)))) add (8.75 mul $x_{2}$)))) add ((-0.9500 mul $x_{3}$) add (-8.1 mul $x_{1}$))) mul $x_{3}$)) add ((-9.74 mul $x_{3}$) add ((((-0.9 mul $x_{3}$) sub (4.45 mul sqrt((-0.373 add (-0.151 mul $x_{3}$))))) add (-54.6 mul $x_{3}$)) sub (-0.758 mul ((85.3000 add (8.74 mul $x_{3}$)))**2))))) $y_{3}$ = (-0.975 add ((-54.4000 mul sqrt((-0.722 add (-9.33 mul $x_{2}$)))) sub (1.45 mul ((66.4 add (-9.65 mul $x_{1}$)))**2))) Symbolic expression of Figure 11 (g) $y_{1}$ = (-7.17 add (0.537 mul $x_{1}$)) $y_{2}$ = (-0.843 add (48.8000 mul $x_{1}$)) $y_{3}$ = (57.3000 add (((-0.449 mul $x_{2}$) add (-1.32 mul $x_{3}$)) add ((-0.9400 mul $x_{4}$) add (0.51 mul $x_{1}$)))) $y_{4}$ = (-0.2040 add (((-6.6000 mul inv((0.88 add (58.1 mul $x_{4}$)))) sub ((-23.0 mul $x_{4}$) add ((-91.0 mul $x_{3}$) sub (-93.6000 mul $x_{2}$)))) sub ((-6.6000 mul $x_{4}$) sub (0.9580 mul (($x_{3}$ mul $x_{3}$) mul ((-0.45 mul $x_{2}$) sub ((((-9.09 mul $x_{4}$) sub ((8.93 mul sqrt(((-26.6 mul $x_{4}$) add (-0.907 mul $x_{1}$)))) add (-6.2 mul $x_{4}$))) sub (-0.078 mul $x_{4}$)) sub (-16.5 mul $x_{2}$))))))))
(a) 1 input channel 1 output channel data example 1
(b) 1 input channel 1 output channel data example 2
(c) 2 input channels 2 output channels data example 1
(d) 2 input channels 2 output channels data example 2
(e) 3 input channels 3 output channels data example 1
(f) 3 input channels 3 output channels data example 2
(g) 4 input channels 4 output channels data example 1
(h) 4 input channels 4 output channels data example 2
Figure 11: Visualization of series from 1 input channel 1 output channel to 4 input channels 4 output channels.
B.2 Composition and Usage of the Series-Symbol Dataset
We set the maximum number of input channels and the maximum number of output channels to 6 and 12 respectively to generate symbolic expressions and series. Each symbolic expression is sampled only once. We generated a total of 40M pairs of series and symbols. The cumulative series length is 50B. The data number of each input channel and output channel in the dataset is shown in Figure 12.
When pre-training SymTime with $S^{2}$ dataset, we start by combining the sampled and generated series and then segmenting them into patches using a sliding window PatchTST ; Time-LLM . The sliding window’s kernel size and step size are both set to 16. Due to the requirement for mask time series modeling (MTM) SimMTM ; HiMTM , there is no overlap between adjacent patches. Given the varying number of input and output channels in the data, the series from the maximum input and output channels can be segmented into up to 288 patches ($18\times 256/16$) video-transformer . For series with fewer than 288 patches, we pad them with zeros to align the length. Next, for symbolic expressions in natural language form BERT ; SNIP ; S2IP-LLM ; Autoformer ; ChatTime , we set a maximum length of 512 characters and perform tokenization. Ultimately, the time series patches and natural language tokens are fed into the time series encoder and the LLM of the Transformer architecture, respectively.
B.3 Statistics Analysis
Setup.
In Section B.2, we provide a detailed introduction to the generation process, composition and usage of the $S^{2}$ dataset SNIP ; Symbolic ; DL4Symbolic . In this section, we first conduct a random sampling analysis of the statistical characteristics of the $S^{2}$ dataset, including stationarity ADF and predictability forecastable ; Timer .
Stationarity.
Stationarity is one of the fundamental properties of time series ARIMA ; ARMA . This attribute ensures that the statistical characteristics of time series data remain consistent across different time points, which is crucial for building effective predictive models and making reliable statistical inferences. To this end, we employ the Augmented Dickey-Fuller (ADF) ADF test to examine the stationarity of the data, thereby determining whether the generated $S^{2}$ dataset is suitable for deep neural networks (DNNs) to learn representations of time series.
Forecastability.
The forecastability of a time series refers to the ability and accuracy to forecast future values based on historical data and statistical models ARIMA_old ; Timer . For certain specific time series and complex systems, such as stock markets, it is often challenging to predict their subsequent developments. Therefore, it is necessary to test whether the $S^{2}$ dataset is non-chaotic and learnable. Forecastability is calculated by subtracting the entropy of the series’ Fourier decomposition as adopted from forecastable and Timer , where a higher forecastability value indicates better predictability. Please note that since the method provided by forecastable is only applicable to multivariate time series, we merge the input channels and output channels together for calculation.
Test Methods and Results.
For the multiple input-output channels presented in the Table 10, we randomly selected 1,000 samples to calculate their average ADF statistics, p-values, and Forecastability metrics. The results indicate that the average p-value from the ADF test across all samples is greater than 0.05, suggesting that the majority of the generated series in the $S^{2}$ dataset are non-stationary time series, posing a challenge in modeling and learning ADF . However, the Forecastability metric, which is greater than 0.3 for all tested samples, indicates that the generated series $Y$ is not produced by a chaotic system and is, overall, predictable.
Figure 12: The number of samples in each part of the $S^{2}$ dataset.
Table 10: Results of the stationarity and forecastability tests for the $S^{2}$ dataset.
inputs outputs ADF p value forecast inputs outputs ADF p value Forecastability
1 1 -12.77 0.0538 0.3155 1 6 -11.48 0.0619 0.3375
2 2 -11.89 0.0568 0.3199 2 6 -11.46 0.0733 0.3218
3 3 -12.40 0.0544 0.3328 3 6 -11.43 0.0625 0.3244
4 4 -11.66 0.0617 0.3491 4 6 -11.53 0.0640 0.3428
5 5 -11.38 0.0628 0.3140 5 6 -12.32 0.0597 0.3284
6 6 -12.43 0.0625 0.3262 6 8 -11.52 0.0555 0.3246
6 10 -11.65 0.0619 0.3287 6 12 -11.66 0.0520 0.3310
B.4 Analysis of Existing Large-scale Datasets for Time Series Pre-training
Table 11: Time-300B time series dataset from Time-MoE Time-MoE .
{footnotesize }
Energy Finance Health Nature Sales Synthetic Transport Web Other Total
# Obs. 15.98B 413.70K 471.04K 279.72B 26.38M 9.22B 2.13B 1.80B 20.32M 309.09B
% 5.17% 5.17% 0.0001% 90.50% 0.008% 2.98% 0.69% 0.58% 0.006% 100%
Table 12: UTSD time series dataset from Timer Timer , where Mise. means Multiple Sources.
{footnotesize }
Energy Environment Health IoT Nature Transport Web Cloud Sales Finance Mise.
# Obs. 16.86B 70.45M 233.M 165M 201B 4.9B 157M 2.15B 198M 0.33M 56.52M
% 7.461% 0.031% 0.103% 0.073% 89% 2.17% 0.07% 0.95% 0.088% 0.00% 0.025%
Table 13: LOTSA time series dataset from Moirai MOIRAI .
{footnotesize }
Energy Transport Climate CloudOps Web Sales Nature Finance Health Total
# Obs. 16.36B 4.90B 4.19B 1.52B 428M 198M 28.55M 24.92M 1.59M 27.65B
% 59.17% 17.73% 15.15% 5.49% 1.55% 0.72% 0.09% 0.10% 0.01% 100%
Table 14: Time series datasets from neural scaling laws Time-Scaling-Laws
{footnotesize }
Transport Climate Energy CloudOps Health Sales Web Total
# Obs. 4.82B 4.73B 2.34B 2.15B 240M 140M 600M 14.46B
% 33.31% 32.71% 16.15% 14.86% 1.61% 0.96% 0.40% 100%
Large-scale datasets are crucial for building foundation models.
Almost all deep learning models today are data-driven, relying on training data Informer ; InceptionTime ; MAE ; Defenders . Therefore, when constructing a pre-trained foundation model for time series, a large-scale and comprehensively representative pre-training dataset is indispensable MOIRAI ; Timer ; BEiT ; CLIP ; SimCLR . The scaling laws of neural networks indicate that the learning effectiveness of deep neural networks is primarily influenced by three factors: the number of model parameters, the size of the training dataset, and the amount of computational resources Neural-Sclaing-Laws ; Time-Scaling-Laws ; scaling-in-llm ; Scaling-Laws-for-Generative . Expanding the scale of the pre-training dataset can effectively improve the model’s generalization capability and performance, and the performance gains from increasing data volume are independent of the model architecture and training methods Time-Scaling-Laws ; Data-driven-weather ; ECG-LLM ; ALERT . Consequently, an increasing number of models are adopting the approach of training larger-scale models on large-scale pre-training datasets to achieve better performance graph-forecasting . This paper surveys the pre-training datasets used by the three current mainstream pre-trained foundation models—Time-MoE Time-MoE , Moirai MOIRAI , and Timer Timer —as well as the datasets utilized in the study of time series scaling laws Time-Scaling-Laws , which are shown in Tables 11, 12, 13 and 14.
Imbalanced domain distribution issues in large-scale time series datasets.
The distribution of data across various domains indicates that the four large-scale time series pre-training datasets all face issues with imbalanced domain data distribution. For instance, domains such as Nature, Energy and Transport have the most datasets Informer , while others like Sales, IoT, Web, Finance and Multiple Sources suffer from extremely low data volumes due to difficulties in data collection or data privacy concerns. According to the scaling laws of neural networks, the imbalance in the pre-training dataset distribution can lead to significant performance biases in in-domain and out-of-domain forecasting tasks for the trained foundation models Time-Scaling-Laws ; graph-forecasting , meaning there is a considerable performance gap between domains with less data and those with more data. To address this, this paper proposes an unrestricted method for generating high-quality time series data to alleviate the scarcity and imbalanced distribution of data in time series analysis domains.
B.5 $S^{2}$ Dataset Statistical Characterization Coverage Experiments
Metric.
To further examine the diversity of the artificially synthesized data in the $S^{2}$ dataset, we conduct a sampling assessment from six dimensions: stationarity, predictability, frequency domain characteristics, complexity, seasonality intensity, and trend characteristics. For each dimension, we select corresponding statistical indicators for dataset evaluation and quantification, as detailed below:
  1. 1.
    Augmented Dickey-Fuller (ADF) Test: Consistent with section B.3, we employ the ADF test to assess the stationarity of time series, using its test statistic as an indicator of time series stationarity ADF ; Timer .
  2. 2.
    Forecastability: Based on forecastable method, we determine whether a time series is chaotic or can be accurately predicted through machine learning models by using Fourier decomposition and entropy Timer . Note that since the method provided by forecastable is only applicable to multivariate time series, we invert the sampled single-channel time series to form a dual-channel series to calculate the indicator.
  3. 3.
    FFT Mean: We utilize the average of the Fourier transform power spectrum to evaluate the frequency domain characteristics of time series. This indicator can be used to measure the overall intensity of time series and assess the energy distribution.
  4. 4.
    Permutation Entropy: This indicator assesses the dynamic complexity of a time series by analyzing its permutation patterns permutation . We set the embedding dimension $m=3$ and time delay $\tau=1$, and calculate its specific value using Shannon Entropy in Equation 8. See permutation for more detailed calculation.
  5. 5.
    Seasonality: We decompose the time series into trend, seasonal and residual components using the Seasonal-Trend Decomposition using LOESS (STL) algorithm seasonality . Then, we calculate the intensity of the seasonal component in the time series according to Equation 9.
  6. 6.
    Mann-Kendall Test: This is a non-parametric statistical method used to detect monotonic trends in time series MK-test . The basic principle is to compare the size relationship between each data point and other data points in the time series. Therefore, this method does not rely on a specific distribution of data and is not affected by outliers. We use the statistical test results of this method as the evaluation indicator, where -1 indicates a downward trend, 1 indicates an upward trend, and 0 indicates no obvious trend.
$\mathrm{Permutation}=-\sum_{j=1}^{K}P_{j}\times\mathrm{ln}P_{j},$ (8)
$\left\{\begin{matrix}Y_{t}=T_{t}+S_{t}+R_{t}\\ \mathrm{Seasonality}=\mathrm{max}\left\{0,1-\frac{\mathrm{Var}(R_{t})}{\mathrm{Var}(S_{t}+R_{t})}\right\}\end{matrix}\right$., (9)
where, $P_{i}$ represents the frequency of the $i$-th permutation model in the permutation entropy, and $K=m!$ is the total number of permutation patterns permutation . $Y_{t}$ represents the original time series, $T_{t}$, $S_{t}$ and $R_{t}$ are the trend, seasonal and residual components decomposed by the STL algorithm seasonality respectively. $\mathrm{Var}(\cdot)$ means calculating the variance of a series.
B.6 Masked Time Series Modeling and Zero-shot Imputation for Representation Learning
Figure 13: Zero-shot time series imputation in S² out-of-domain data.
(a) ETTm1
(b) ETTm2
(c) ETTh1
(d) ETTh2
(e) Electricity
(f) Weather
Figure 14: Zero-shot time series imputation in real world time series dataset in ETTm1, ETTm2, ETTh1, ETTh2 Informer , Electricity ECL and Weather weather .
Setup.
Since we incorporate MTM loss in the pre-training process of SymTime, in this section, we assess the specific learning effects of the time series encoder in SymTime through masked modeling PatchTST ; SimMTM ; TimeSiam ; TimeMixer . We test the model’s performance using both pre-trained synthetic data not in the S² dataset and real datasets from time series imputation tasks SNIP ; Symbolic ; DL4Symbolic . As SymTime adds masks in units of patches of length 16 during pre-training, we also add masks in the form of 16-length patches. The reconstruction effect of the masked parts by the time series encoder is shown in Figure 13 and 14.
$S^{2}$ Dataset Out-of-Domain Data.
In Figure 13, we generate new data using the method from the $S^{2}$ dataset and add masks to test the reconstruction ability of the time series encoder SNIP ; Symbolic ; DL4Symbolic . The gray sections represent the masked segments, while blue and orange represent the original and reconstructed series, respectively. We input time series outside the gray parts in patches and have the model reconstruct the gray sections based on the remaining information. Since we only calculate the MTM loss on the masked parts SimMTM ; TimeSiam , the visible reconstruction does not overlap with the original input series Timer ; TimeFM ; Moment . From the Figure 13, it can be observed that the time series encoder in SymTime performs well in fitting the fluctuations and trends of time series, demonstrating that our encoder successfully learned the fundamental representations of time series during pre-training MOIRAI ; SAMformer ; Discovering .
Real-world Time Series Data.
In Figure 14, we conduct representation learning tests on 6 real datasets: ETTm1, ETTm2, ETTh1, ETTh2 Informer , Electricity ECL , and Weather weather . Since no real data are used for model pre-training, these datasets are also considered as out of domain data. We similarly add masks in patch units (gray sections). It can be observed that the time series encoder in SymTime also performs well in zero-shot reconstruction on real-world data SimMTM ; MaskTime .
B.7 Time Complexity Analysis of $S^{2}$ Data Generation Mechanism
We define the specific symbol and its explanation in Table 9. Then, we use the divide-and-conquer approach to anaylze the complexity of the $S^{2}$ data generation mechanism.
  1. 1.
    Symbolic Expression Generation: We construct symbolic expressions using a tree structure as a medium. When we have $b$ binary operators, we further insert $(b+1)$ leaf nodes (the process from (a) to (b) in Figure 3 in our paper). Therefore, after inserting $u$ unary operators (Figure 3 (c)), the total number of nodes in the tree is $n=2b+u+1$. Because there are many ways to construct a tree, we consider the time complexity of constructing a balanced tree. Therefore, for $N$ symbols constructed, the specific complexity of this process is $\mathcal{O}(N\times n\mathrm{log}n)$.
  2. 2.
    Sampling series generation: When we want to generate a sampling time series with $M$ channels, each channel has a probability of $P$ to be sampled using a mixture distribution and a probability of (1P)(1-P) to be sampled using an ARMA model. When the sampling length of the series is $L$, the complexity of generating $k$ mixture distribution and ARMA ($p$, $q$) series is $O(kL)$ and $O(L(p+q))$. Therefore, the time complexity of this process can be quantified as $\mathcal{O}\left(ML\times[Pk+(1-P)(p+q)]\right)$.
  3. 3.
    Sampling through symbolic expressions and series: We simplify the specific operational details of this process and only consider the time complexity of operations with variables. For a series of length L, we have $N$ symbolic expressions to be sampled, and each symbol has an average of $\frac{M+1}{2}$ variables (Each symbolic expression may contain any number of variables from 1 to M, so here we take $\frac{M+1}{2}=\frac{(1+2+\cdots+M)}{M}$ as the average probability). Then the process can be quantified as $\mathcal{O}(N\cdot\frac{M+1}{2}\cdot L)$.
To sum up, the symbolic expressions we construct and the parameters used in the sampling process are typically smaller than the length of the time series $L$. Therefore, we ignore the symbolic expression generation process and consider only the two processes of generating the sampling series and sampling using the symbolic expression. Since the number of channels, $M$ and $N$, as well as $k$, $p$ and $q$ are all smaller than $L$, we can intuitively assume that the time complexity of the $S^{2}$ data generation mechanism is linearly related to the length of the series $L$.
B.8 The Selection of the Unary Operators
From the perspective of constructing symbolic expressions in the current $S^{2}$ data generation mechanism, binary operators primarily serve to connect multiple variables, while unary operators can increase the diversity of numerical values through specific operations. However, we choose not to use all symbolic and linear operations, but only use the unary operators $\{\mathrm{inv},\mathrm{abs}$, $\mathrm{pow2}$, $\mathrm{pow3},\mathrm{sqrt,sin,cos,tan,arctan,log,exp}\}$ for generation.
Although ignoring certain mathematical symbols will reduce the diversity of symbolic expressions, we have found in numerous experiments and tests that differential $\frac{\mathrm{d}y}{\mathrm{d}x}$, integral $\int$, power operations $x^{n}$, and exponential operations with various bases $n^{x}$ will seriously affect our data generation to a certain extent. The specific reasons are as follows:
  1. 1.
    Value explosion: To maintain quality, we cap large magnitudes (Section 3.1). Integration, exponential and high-order powers readily cause overflow, so they were dropped; exp alone is retained for diversity.
  2. 2.
    Numerical differentiation: Symbolic differentiation introduces truncation/round-off trade-offs. Combined with reciprocal and absolute-value operators, functions such as $|x|$ or $\sin\left(\frac{1}{x}\right)$ because non-smooth or high-frequency vibration at $x=0$, breaking differentiation.
  3. 3.
    Numerical integration: Randomly built expression trees often yield integrands with singularities (e.g., $\int_{0}^{1}\frac{1}{\sqrt{x}}\mathrm{d}x$) or force costly oscillatory integrals (e.g., $\int_{0}^{100}sin(100x)\mathrm{d}x$); interval selection is non-trivial.
  4. 4.
    Symbolic cost: Differentiation and integration are slow and can trigger exponential memory growth. Many elementary functions lack closed-form antiderivatives (e.g., $\int\exp(-x^{2})\mathrm{d}x$).
Considering factors like numerical stability, symbolic complexity, computational efficiency, and sampling success rate, we selectively omitted some symbolic operations. Nevertheless, the data generation framework proposed is essentially a complete theory. It already incorporates the vast majority of symbolic operations, and new operators or user-defined symbolic operations can be easily added. The omission of some operations due to the above factors does not undermine the validity of this framework.
Figure 15: The time complexity analysis of $S^{2}$ Generation
To further demonstrate that the time complexity of the $S^{2}$ data generation mechanism is linearly related $\mathcal{O}(L)$ to the length of the generated time series, we start with a time series of length 16 and generate it every 16 lengths until 512. We switch to a different random seed for each generation and repeat the experiment 1280 times. The average time of each data generation is shown in Figure 15. The linear fit line of the result shows that the time complexity of data generation is linearly related to the sequence length when sampling failure is not considered.
B.9 The Limitations of $S^{2}$ Generation Mechanism
As outlined in the abstract and introduction, to address the scarcity of training data for time series foundation models, this work proposes a dual-modal data generation mechanism grounded in complex dynamical systems theory (detailed in Section 3.1). This mechanism enables comprehensive coverage of time series representation spaces through unrestricted, high-quality generation. However, generating symbolic expressions (complex systems) via randomized binary tree algorithms occasionally results in oversized trees, leading to overly intricate symbolic systems. This issue narrows the domain of symbolic functions $f(\cdot)$, hinders sampling of stimulus-driven time series $X$, and reduces sampling efficiency. Notably, differential and integral operations—complex linear transformations—severely degrade sampling speed. Consequently, these operations are excluded from the current $S^{2}$ dataset generation. Future work will explore integrating ordinary and partial differential equations into the $S^{2}$ framework to enrich symbolic expression diversity (complex systems) and further enhance the representational capacity of generated time series data.
Appendix C Implementation Details
In this section, we first provide a detailed introduction to the datasets and evaluation metrics used for the five TSA tasks. Subsequently, we elaborate on the training details of our experiments, including how we pre-trained SymTime on the $S^{2}$ dataset and how we fine-tuned it on downstream task datasets. All experiments and deep neural networks training are implemented in PyTorch on 8 NVIDIA A6000 48GB GPU.
C.1 Downstream Tasks Datasets Details
Table 15: Dataset descriptions. The dataset size is organized in (Train, Validation, Test).
{footnotesize }
Tasks Dataset Dim Series Length Dataset Size
Information (Frequency)
ETTm1, ETTm2 7
{96, 192, 336, 720}
(34465, 11521, 11521)
Electricity (15 mins)
ETTh1, ETTh2 7
{96, 192, 336, 720}
(8545, 2881, 2881)
Electricity (15 mins)
Forecasting Electricity 321
{96, 192, 336, 720}
(18317, 2633, 5261)
Electricity (Hourly)
(Long-term) Traffic 862
{96, 192, 336, 720}
(12185, 1757, 3509)
Transportation (Hourly)
Weather 21
{96, 192, 336, 720}
(36792, 5271, 10540)
Weather (10 mins)
Exchange 8
{96, 192, 336, 720}
(5120, 665, 1422)
Exchange rate (Daily)
M4-Yearly 1 6 (23000, 0, 23000)
Demographic
M4-Quarterly 1 8 (24000, 0, 24000)
Finance
Forecasting M4-Monthly 1 18 (48000, 0, 48000)
Industry
(short-term) M4-Weakly 1 13 (359, 0, 359)
Macro
M4-Daily 1 14 (4227, 0, 4227)
Micro
M4-Hourly 1 48 (414, 0, 414)
Other
Imputation ETTm1, ETTm2 7 96 (34465, 11521, 11521)
Electricity (15 mins)
ETTh1, ETTh2 7 96 (8545, 2881, 2881)
Electricity (15 mins)
Electricity 321 96 (18317, 2633, 5261)
Electricity (15 mins)
Weather 21 96 (36792, 5271, 10540)
Weather (10 mins)
EthanolConcentration
3 1751 (261, 0, 263)
Alcohol Industry
FaceDetection
144 62 (5890, 0, 3524)
Face (250Hz)
Handwriting
3 152 (150, 0, 850)
Handwriting
Heartbeat
61 405 (204, 0, 205)
Heart Beat
Classification
JapaneseVowels
12 29 (270, 0, 370)
Voice
PEMS-SF
963 144 (267, 0, 173)
Transportation (Daily)
SelfRegulationSCP1
6 896 (268, 0, 293)
Health (256Hz)
SelfRegulationSCP2
7 1152 (200, 0, 180)
Health (256Hz)
SpokenArabicDigits
13 93 (6599, 0, 2199)
Voice (11025Hz)
UWaveGestureLibrary
3 315 (120, 0, 320)
Gesture
SMD 38 100 (566724, 141681, 708420)
Server Machine
Anomaly MSL 55 100 (44653, 11664, 73729)
Spacecraft
Detection SMAP 25 100 (108146, 27037, 427617)
Spacecraft
SWaT 51 100 (396000, 99000, 449919)
Infrastructure
PSM 25 100 (105984, 26497, 87841)
Server Machine
We conduct experiments using the TimesNet benchmark TimesNet , with a detailed description of the dataset provided in Table 15. Specifically, we utilize 8 datasets including ETTh1, ETTh2, ETTm1, ETTm2 Informer , Electricity ECL , Traffic traffic , Weather weather , and Exchange LSTNet to conduct long-term time series forecasting experiments. Our model, SymTime, employ input series of lookback lengths 96 and 512, with forecast horizons of 96, 192, 336, and 720. For short-term forecasting experiments, we employ the M4 benchmark dataset, predicting data of various frequencies M4team2018dataset . In the time series imputation task, we test on 6 datasets—ETTh1, ETTh2, ETTm1, ETTm2 Informer , Electricity ECL , and Weather weather —with mask rates of 12.5%, 25%, 37.5%, and 50%. For time series classification, we utilize ten UEA multivariate time series classification benchmark datasets UEA . For anomaly detection in time series, we experiment with five datasets: SMD SMD , MSL MSL , SMAP MSL , SWaT SWaT , and PSM PSM .
C.2 Metrics
We assess the five TSA tasks using various metrics. For long-term forecasting and imputation tasks, we employ mean squared error (MSE) and mean absolute error (MAE). For short-term forecasting, we utilize symmetric mean absolute percentage error (SMAPE), mean absolute scaled Error (MASE), and overall weighted average (OWA), with OWA being a metric unique to the M4 competition. For time series classification tasks, we use classification accuracy as the metric. For anomaly detection tasks, we adopt precision, recall, and F1-score as our evaluation metrics. The calculations for these metrics are as follows.
$\mathrm{MSE}=\sum_{i=1}^{n}\left(y_{i}-\hat{y}_{i}\right)^{2},$ (10)
$\mathrm{MAE}=\sum_{i=1}^{n}\left|y_{i}-\hat{y}_{i}\right|,$ (11)
$\mathrm{SMAPE}=\frac{200}{T}\sum_{i=1}^{T}\frac{\left|\mathbf{X}_{i}-\hat{\mathbf{Y}_{i}}\right|}{\left|\mathbf{X}_{i}\right|+\left|\hat{\mathbf{Y}}_{i}\right|},$ (12)
$\mathrm{MAPE}=\frac{100}{T}\sum_{i=1}^{T}\frac{\left|\mathbf{X}_{i}-\hat{\mathbf{Y}_{i}}\right|}{\left|\mathbf{X}_{i}\right|},$ (13)
$\mathrm{MASE}=\frac{1}{T}\sum_{i=1}^{T}\frac{\left|\mathbf{X}_{i}-\hat{\mathbf{Y}_{i}}\right|}{\frac{1}{T-q}\sum_{j=q+1}^{T}\left|\mathbf{X}_{j}-\mathbf{X}_{j-q}\right|},$ (14)
$\mathrm{OWA}=\frac{1}{2}\left[\frac{\mathrm{SMAPE}}{\mathrm{SMAPE}_{\text{Naïve2}}}+\frac{\mathrm{MASE}}{\mathrm{MASE}_{\text{Naïve2}}}\right],$ (15)
where, $y_{i}$ is the ground true value, $\hat{y}_{i}$ is the model prediction, $q$ is the peridoicity of the time series data. $\mathbf{X},\hat{\mathbf{Y}}\in\mathbb{R}^{T\times C}$ are the ground truth and prediction results of the future with $T$ time points and $C$ dimensions. $\mathbf{X}_{i}$ means the $i$-th future time point.
C.3 Pre-training
Model Architecture.
The model architecture of the time series and symbolic encoders in SymTime are shown in Table 16.
Table 16: The model architecture of the time series and symbolic encoders in SymTime.
Encoder Layers $d_{\mathrm{model}}$ $d_{\mathrm{ff}}$ Heads Params
Time 6 512 2048 8 19M
Symbol 6 786 3072 12 67M
Model Hyper-parameter.
The parameter configurations for the time series encoder and symbol encoder in SymTime are shown in Table 16. During model pre-training, we primarily set three hyperparameters: (1) the masking ratio of time series patches, (2) the masking ratio for natural language symbols, and (3) the proportion factor $\alpha$ used to balance pseudo-targets in momentum distillation. Based on the masked time series modeling pre-training experimental configuration of PatchTST PatchTST and SimMTM SimMTM , we set the masking ratio for time series to 40%. Following the experimental configuration of BERT in masked language modeling BERT ; DistilBERT , we set the masking ratio for symbolic data to 15%. Based on the experimental configuration of momentum distillation in ALBEF ALBEF ; momentum-distillation ; MoCo_v1 , we set $\alpha$ to 0.6.
Training Configurations.
During the pre-training of SymTime, we employ AdamW Adam ; AdamW as the optimizer with the defult hyperparameter configuration for ($\beta_{1}$, $\beta_{2}$) as (0.9, 0.999). Then, we utilize the OneCycle policy to dynamically adjust the learning rate. We set the warmup epochs to 10, during which the learning rate gradually grows up to an initial value of $5\times 10^{-5}$, and then adjust it dynamically using a cosine annealing schedule, with the minimum learning rate set at $1\times 10^{-7}$. We conduct pre-training using data parallelism on a hardware setup consisting of 8 NVIDIA RTX A6000 GPUs with 48GB of memory each. We set the batch size to 128 and trained for a total of 85 epochs. Unlike SNIP SNIP , we do not generate data on-the-fly during training for pre-training. Instead, we prepare the data in advance and then load it into the device for pre-training. Due to the large size of our generat $S^{2}$ dataset, we load data into the GPU in batches during each epoch for pre-training.
C.4 Fine-tuning
For the five major tasks in TSA, we conduct downstream task fine-tuning experiments using the configurations in Table 17. For all downstream task fine-tuning experiments, we employ the Adam optimizer Adam ; AdamW with hyperparameters $(\beta_{1},\beta_{2})$ set to $(0.9,0.999)$. The LR in the table represents the initial learning rate and we utilize the dynamic learning rate adjustment strategy from TimesNet TimesNet .
Table 17: Experiment configuration of SymTime fine-tuning.
Tasks / Configurations Model Parameter Training Configurations
$d_{\mathrm{model}}$ $d_{\mathrm{ff}}$ Layers LR Loss Batch Size Epochs
Long-term Forecasting 512 2048 3, 6 $10^{-4}-5\times 10^{-4}$ MSE 4-64 20
Short-term Forecasting 2, 3 $10^{-4}-2\times 10^{-4}$ SMAPE 8-32 16
Classification 1-6 $10^{-4}-5\times 10^{-3}$ Cross Entropy 4-64 64
Imputation 2, 3, 6 $10^{-4}-5\times 10^{-4}$ MSE 4-64 32
Anomaly Detection 3, 6 $10^{-4}-5\times 10^{-4}$ MSE 4-64 12
C.5 Ablation Experiments Details
Ablation study on pre-training strategies and objectives.
To further verify the effectiveness of our series-symbol pre-training strategy and objectives, we establish 8 distinct ablation experiment groups and a control group. The specific configurations of these 8 ablation experiment groups are as follows.
  1. 1.
    Freeze: All parameters in the pre-trained time series encoder are frozen, with only the linear projection layer for outputting prediction results fine-tuned.
  2. 2.
    Read-Data: Since real time series data does not have matching symbolic expression information, we temporarily discarded the symbolic encoder and momentum model in this ablation experiment and only used real time series data for pre-training using the MTM method.
  3. 3.
    w/o Pretrain: No series-symbol pre-training is conducted; the time series encoder with initialized parameters is used for downstream task experiments.
  4. 4.
    w/o MTM: The masked time series modeling (MTM) is removed from the pre-training objectives.
  5. 5.
    w/o MLM: The masked language modeling (MLM) is removed from the pre-training objectives.
  6. 6.
    w/o T2S: The contrastive loss from time series to symbols is removed from the pre-training objectives.
  7. 7.
    w/o S2T: The contrastive loss from symbols to time series is removed from the pre-training objectives.
  8. 8.
    w/o Symbol: Only time series data from the $S^{2}$ dataset are used to pre-train the time series encoder via MTM, disregarding the correspondence with symbols.
  9. 9.
    w/o Distill: The contrastive loss in pre-training does not use the pseudo objective of momentum distillation.
C.6 Ablation Experiments on Short-term Forecasting
Setup.
We adopt the same experimental setup as in Section 4.4 to conduct ablation studies on short-term time series forecasting tasks. We first select the Yearly and Monthly sub-datasets from the M4 benchmark dataset M4team2018dataset to perform ablation experiments on SymTime’s pre-training objectives. We choose SMAPE as the evaluation metric and the average results with error bars are shown in Figure 16.
Figure 16: Ablation study on short-term forecasting task.
Results.
Figure 16 (a) indicates that SymTime’s performance drops sharply when the backbone encoder is frozen and no pre-training is conducted. When some pre-training objectives are removed, the model’s performance in short-term time series forecasting also declines, but the sensitivity of performance degradation is not as pronounced as in long-term forecasting experiments. Figure 16 (a) shows that as the size of the pre-training dataset increases, SymTime’s performance on the Quarterly dataset improves significantly.
C.7 The Ablation of Backbone in SymTime
Setup.
SymTime is composed of two encoders with Transformer architectures Transformer . The time series encoder is composed of a multi-layer Transformer encoder architecture. The symbolic expression encoder is pre-trained with a large language model. We conducted an ablation experiment on the SymTime model architecture by changing the encoder structure through control variables. Specifically, we adjusted the number of parameters ($d_{\mathrm{model}}$ and $d_{\mathrm{ff}}$) of the time series encoder and the type of pre-trained LLM used by the symbolic encoder to set different control groups:
  • SymTime: The original model architecture in Table 16 uses the DistilBERT DistilBERT .
  • SymTimesmall: Change the time series encoder to a 3-layer Transformer model with $d_{\mathrm{model}}=386$ and $d_{\mathrm{ff}}=1536$.
  • SymTimelarge: Change the time series encoder to a 6-layer Transformer model with $d_{\mathrm{model}}=768$ and $d_{\mathrm{ff}}=3072$.
  • BERT-base110M: Replace the pre-trained LLM with BERT-base110M BERT .
  • BERT-large340M: Replace the pre-trained LLM with BERT-large340M BERT .
  • GPT2-small124M: Replace the pre-trained LLM with GPT2-small124M GPT-2 .
  • GPT2-medium335M: Replace the pre-trained LLM with GPT2-medium335M GPT-2 .
Figure 17: The ablation results of backbone in SymTime. We choose to conduct experimental verification on four ETT datasets Informer for long-term time series prediction.
Results.
As shown in Figure 17, changing the backbone of SymTime does not significantly affect the experimental results in long-term time series forecasting tasks. Ablation experiments on pre-training objectives (Equation 7) reveal that the performance gains achieved during pre-training are primarily due to the pre-training paradigm of masked modeling of time series and symbolic expressions and contrastive learning. This pre-training approach is independent of the model backbone. Therefore, if the basic pre-training requirements are met (pre-training loss can be successfully optimized), a more lightweight model can be used for fine-tuning on downstream tasks.
C.8 The Impact and Ablation of Pre-interpolation on Time Series Imputation Task
SymTime adds masks randomly at the patch level during pre-training for time series reconstruction. While in the imputation task, masks are added randomly at the data point level. Additionally, high masking rates may disrupt the original trend and periodic features of the time series. Therefore, we use Peri-midFormer’s method to apply per-interpolation to the masked time series to restore the disrupted periodic features Peri-midFormer ; TimesNet ; TimeMixer ; TimeMixer++ ; SIN ; CauDiTS . It is important to note that this method is general and independent of deep learning models. The use of this method aims to further enhance the potential of deep learning models. To further verify the effectiveness and impact of the per-interpolation method, we conduct experiments on the ECL time series imputation dataset, with results shown in Table 18. Taking the ECL time series dataset 0.5 mask ratio as an example, the effect of pre-interpolation is shown in Figure 18. We perform experiments with masking rate of $\{0.125,0.25,0.375,0.50\}$ and compare models such as Peri-midFormer Peri-midFormer , TimesNet TimesNet , PatchTST PatchTST , DLinear DLinear and Pyraformer Pyraformer . Per-interpolation represents the experimental results obtained using only linear interpolation. For a missing time series $x_{t}$ at time $t$, the method can be described as:
$x_{t}=\left\{\begin{matrix}\frac{x_{t-1}+x_{t+1}}{2},&if\ (x_{t-1})\neq\mathrm{None}\ \&\ (x_{t+1}\neq\mathrm{None})\\ x_{t+1},&if\ (x_{t-1})=\mathrm{None}\ \&\ (x_{t+1}\neq\mathrm{None})\\ x_{t-1},&if\ (x_{t-1})\neq\mathrm{None}\ \&\ (x_{t+1}=\mathrm{None})\end{matrix}\right$., (16)
where, $x_{t-1}$ and $x_{t+1}$ represent the values at the previous and next time points, respectively, while None indicates a missing value. The results in Table 18 show that this method significantly improves the performance of all models in a model-independent manner.
(a) Original data
(b) Data with 50% missing values
(c) Per-interpolated data
Figure 18: Visualization of original data, data with 50% missing values and pre-interpolated data of ECL dataset.
Table 18: Ablation Experiments of pre-interpolation in inputation task on ECL dataset. The per-interpolation results for Peri-midFormer, TimesNet, PatchTST, DLinear and Pyraformer are copied from Peri-midFormer .
Methods Metric w/o per-interpolation with per-interpolation
0.125 0.25 0.375 0.5 0.125 0.25 0.375 0.5
Per-interpolation MSE - - - - 0.086 0.110 0.149 0.206
MAE - - - - 0.188 0.213 0.251 0.301
SymTime (Ours) MSE 0.050 0.064 0.074 0.092 0.037 0.047 0.060 0.075
MAE 0.145 0.169 0.181 0.206 0.122 0.139 0.160 0.181
Peri-midFormer Peri-midFormer MSE 0.073 0.092 0.107 0.122 0.047 0.053 0.067 0.085
MAE 0.187 0.214 0.231 0.248 0.140 0.162 0.179 0.195
TimesNet TimesNet MSE 0.088 0.092 0.096 0.102 0.081 0.083 0.086 0.091
MAE 0.203 0.208 0.214 0.221 0.196 0.198 0.201 0.207
PatchTST PatchTST MSE 0.061 0.072 0.082 0.097 0.050 0.059 0.070 0.087
MAE 0.170 0.185 0.198 0.216 0.148 0.164 0.181 0.202
DLinear DLinear MSE 0.084 0.113 0.141 0.173 0.050 0.062 0.789 0.105
MAE 0.206 0.243 0.273 0.303 0.144 0.164 0.189 0.225
Pyraformer Pyraformer MSE 0.297 0.294 0.296 0.299 0.165 0.165 0.171 0.173
MAE 0.383 0.380 0.381 0.383 0.290 0.291 0.293 0.295
Appendix D Related Work
D.1 Time Series Foundation Models
In CV and NLP CLIP , PTFMs have been demonstrated to adapt to a variety of downstream tasks after fine-tuning on specific datasets, exhibiting excellent generalization and scalability. Inspired by this, recent years have seen significant progress in PTFMs for TSA Transformer-in-TSA ; KDD-Survey ; TimeDiT , with the emergence of various pre-training methods. MOIRAI, through MTM and reconstruction, has been pre-trained on large datasets (27B), yielding a universal forecasting model with significant zero-shot advantages MOIRAI . Timer, after generative pre-training on large datasets (1B), has performed well in forecasting Timer . TimeGPT trained a encoder-decoder Transformer with 100B data TimeGPT . COMET, using multi-level contrastive learning on a large ECG dataset, has obtained a medical time series PTFMs with few-shot advantages COMET .
As discussed in Appendix Section B.4, these baseline models still face challenges related to data scarcity and data imbalance. In the next section, we introduce the proposed data generation mechanism and the corresponding dual-modality foundation model designed to address these issues.
D.2 Deep Learning and Symbolic Regression
The central thesis of this paper is to regard time series as representations of complex dynamical systems AI-Feynman . Traditionally, complex systems are modeled by observing time series utilizing ODE and PDE DE4ModelingCS . With the advancement of machine learning, symbolic regression (SR) Exhaustive , as a supervised learning method, can discover hidden mathematical expressions from numerical series. Although genetic algorithms (GAs) are the mainstream approach for SR GA4SR1 ; GA4SR2 , deep learning-based methods have also made significant progress. Symbolic constructed an end-to-end SR model using Transformers, while SNIP built a large-scale pre-trained model through contrastive learning on symbolic expressions and numerical observations SNIP . Both methods treat symbolic expressions as nature language and use deep neural networks to learn their features. Therefore, this paper employs a pre-trained LLM as a symbol encoder to learn the features of symbolic expressions and jointly trains a time series foundation model imbued with semantic information through contrastive learning CauDiTS .
D.3 Time Series Forecasting Models Based on Synthetic Data
Unlike the representation pre-training conducted on the large synthetic $S^{2}$ dataset in this paper, previous TSA models trained on synthetic data were mainly based on Prior-data Fitted Networks (PFN) TDBI ; PFN . This model learns prior distributions from synthetic data using Bayesian methods, enabling zero-shot inference. ForecastPFN generated a large number of synthetic time series by separately modeling the seasonal trend, global trend and noise based on given constraint expressions ForecastPFN . Although PFN trained in this way offered certain zero-shot and few-shot advantages, this approach was limited to generating time series through sampling fixed expressions and performing linear combinations. In contrast, the $S^{2}$ data generation mechanism proposed in this paper can sample an infinite variety of symbolic expressions SNIP ; Symbolic ; DL4Symbolic ; CauDiTS . TimePFN constructed synthetic datasets by filtering real time series with linear and periodic convolution kernels, training PFN for zero-shot inference TimePFN . However, this method depends on real-world time series for filtering and linear transformations between channels. Compared to the $S^{2}$ data generation mechanism, it can not create large-scale and fully representative synthetic datasets for model pre-training.
Appendix E Visualization
E.1 Long-term Time Series Forecasting with 96 Prediction on ETTh1 (Figure 19) and ECL (Figure 20)
E.2 Short-term Time Series Forecasting on M4 Weekly (Figure 21) and Monthly (Figure 22)
E.3 Time Series Imputation with 50% mask rate on ETTh1 (Figure 23) and ETTm1 (Figure 24)
Appendix F Full Results
For the five downstream TSA tasks results, we use (1) Peri-midFormer for Peri-midFormer Peri-midFormer , (2) uni2ts for Moirai MOIRAI , (3) Large-Time-Series-Model for Timer Timer , (4) Time-LLM for Time-LLM Time-LLM , (5) TSLANet for TSLANet TSLANet , (6) S2IP-LLM for S2IP-LLM S2IP-LLM , (7) NeurIPS²023-One-Fits-All for GPT4TS GPT4TS , (8) UniTS for UniTS UniTS , (9) moment for Moment Moment , (10) FilterNet for FilterNet FilterNet , (11) RTSF for RLinear RLinear , and (12) Time-Series-Library for other models, such as TimesNet TimesNet , PatchTST PatchTST , TimeMixer TimeMixer , iTransformer iTransformer , DLinear DLinear , Autoformer Autoformer and Informer Informer , TimeXer TimeXer , Chronos-forecasting for Chronos Chronos . To ensure a fair comparison, we use the original experimental configuration in the project scripts.
F.1 Time Series Long-term Forecasting with 96 look-back windows (Table 19, Table 20 and Table 21)
F.2 Time Series Long-term Forecasting with 336 look-back windows (Table 22)
F.3 Time Series Long-term Forecasting with 512 look-back windows (Table 23)
F.4 Time Series Short-term Forecasting (Table 24 and Table 25)
F.5 Time Series Classification (Table 26 and Table 27)
F.6 Time Series Imputation (Table 28 and Table 29)
F.7 Time Series Anomaly Detection (Table 30)
F.8 Pre-training and Fine-tuning Results of Long-term Forecasting (Table 31)
F.9 Pre-training and Fine-tuning Results of Short-term Forecasting (Table 32)
F.10 Pre-training and Fine-tuning Results of classification (Table 33)
F.11 Pre-training and Fine-tuning Results of Imputation (Table 34)
F.12 Pre-training and Fine-tuning Results of Anomaly Detection (Table 35)
Appendix G Impact Statement
The potential value of this work lies in its ability to mitigate fundamental challenges in TSA, such as the lack of sufficient labeled data and the issue of imbalanced datasets. By generating rich, diverse, and high-quality synthetic data, our approach not only addresses these issues but also opens new avenues for improving model generalization across a wide range of applications. Furthermore, the dual-modality framework, which combines time series data with symbolic semantics, introduces a novel way of enriching the representation power of models, allowing them to better understand complex temporal dynamics and their underlying patterns.
We foresee that pre-training models on synthetic datasets, especially those that combine structured symbolic information with time series data, will become a key development trend in the TSA field. This could pave the way for more robust and scalable solutions in a variety of domains, including finance, healthcare, and climate modeling, where time series data is abundant, but labeled data is often scarce or hard to obtain.
(a) SymTime
(b) PatchTST
(c) iTransformer
(d) TimesNet
(e) DLinear
(f) Autoformer
Figure 19: Visualization of long-term forecasting with 96 prediction length of ETTh1 dataset.
(a) SymTime
(b) PatchTST
(c) iTransformer
(d) TimesNet
(e) DLinear
(f) Autoformer
Figure 20: Visualization of long-term forecasting with 96 prediction length of Electricity dataset.
(a) SymTime
(b) PatchTST
(c) iTransformer
(d) TimesNet
(e) DLinear
(f) Autoformer
Figure 21: Visualization of time series short-term forecasting in M4 dataset Weekly.
(a) SymTime
(b) PatchTST
(c) iTransformer
(d) TimesNet
(e) DLinear
(f) Autoformer
Figure 22: Visualization of time series short-term forecasting in M4 dataset Monthly.
(a) SymTime
(b) PatchTST
(c) iTransformer
(d) TimesNet
(e) DLinear
(f) Autoformer
Figure 23: Visualization of time series imputation with 50% mask rate of ETTh1 dataset.
(a) SymTime
(b) PatchTST
(c) iTransformer
(d) TimesNet
(e) DLinear
(f) Autoformer
Figure 24: Visualization of time series imputation with 50% mask rate of ETTm1 dataset.
Table 19: Full results for the long-term forecasting task compared with Peri-midFormer Peri-midFormer , Moirai MOIRAI , Timer Timer , Moment Moment , Time-LLM Time-LLM , TSLANet TSLANet , $S^{2}$IP-LLM S2IP-LLM and GPT4TS GPT4TS . (* means former, T-LLM is Time-LLM, $S$-LLM is $S^{2}$IP-LLM.) To ensure fairness in the comparison, we set the look-back window length of all models to 96. Since the Timer and Moirai need to input a longer series to build a token, their windows are 672. $S^{2}$IP-LLM has a gradient explosion when the window is 96, so its look-back window is 512. The standard deviation is within 0.5%. Red: best, Blue: second best.
{footnotesize }
SymTime Peri-mid* Moirai Timer Moment T-LLM TSLANet $S$-LLM GPT4TS
Methods (Ours) Peri-midFormer MOIRAI Timer Moment Time-LLM TSLANet S2IP-LLM GPT4TS
Metrics MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
96 0.318 0.353 0.334 0.370 0.311 0.358 0.315 0.354 0.305 0.353 0.304 0.359 0.321 0.362 0.325 0.371 0.293 0.362
192 0.362 0.380 0.382 0.391 0.381 0.402 0.369 0.378 0.369 0.386 0.368 0.396 0.361 0.383 0.361 0.397 0.374 0.392
336 0.386 0.402 0.417 0.418 0.436 0.432 0.425 0.428 0.392 0.404 0.383 0.393 0.383 0.404 0.385 0.403 0.389 0.404
720 0.419 0.423 0.501 0.461 0.466 0.476 0.442 0.447 0.425 0.436 0.420 0.429 0.445 0.437 0.426 0.446 0.421 0.423
ETTm1
Avg 0.371 0.390 0.409 0.410 0.398 0.417 0.388 0.402 0.373 0.395 0.369 0.394 0.377 0.397 0.374 0.404 0.369 0.395
96 0.174 0.257 0.174 0.255 0.179 0.267 0.168 0.254 0.170 0.264 0.177 0.269 0.179 0.261 0.174 0.263 0.171 0.265
192 0.238 0.299 0.249 0.305 0.244 0.311 0.429 0.425 0.285 0.294 0.239 0.305 0.243 0.303 0.232 0.306 0.226 0.304
336 0.295 0.337 0.319 0.349 0.335 0.371 0.476 0.457 0.275 0.329 0.301 0.340 0.308 0.345 0.300 0.344 0.288 0.345
720 0.390 0.392 0.418 0.405 0.425 0.444 0.545 0.497 0.383 0.397 0.382 0.381 0.403 0.401 0.359 0.386 0.372 0.398
ETTm2
Avg 0.274 0.321 0.290 0.328 0.296 0.348 0.405 0.408 0.278 0.321 0.275 0.324 0.283 0.327 0.266 0.325 0.264 0.328
96 0.376 0.400 0.382 0.403 0.369 0.408 0.374 0.404 0.385 0.402 0.386 0.395 0.387 0.405 0.380 0.403 0.388 0.399
192 0.428 0.431 0.436 0.435 0.441 0.450 0.430 0.438 0.449 0.450 0.421 0.424 0.448 0.436 0.410 0.427 0.425 0.429
336 0.463 0.456 0.492 0.455 0.469 0.469 0.458 0.453 0.455 0.472 0.438 0.450 0.451 0.437 0.426 0.442 0.444 0.455
720 0.450 0.458 0.508 0.490 0.486 0.490 0.475 0.480 0.480 0.503 0.506 0.510 0.505 0.485 0.610 0.543 0.479 0.477
ETTh1
Avg 0.430 0.436 0.455 0.446 0.441 0.454 0.434 0.444 0.442 0.457 0.438 0.445 0.448 0.441 0.456 0.454 0.434 0.440
96 0.293 0.347 0.312 0.358 0.288 0.350 0.315 0.360 0.285 0.343 0.307 0.369 0.289 0.345 0.292 0.353 0.292 0.351
192 0.364 0.397 0.388 0.403 0.390 0.426 0.411 0.423 0.368 0.403 0.349 0.384 0.362 0.391 0.355 0.388 0.351 0.394
336 0.385 0.423 0.443 0.443 0.441 0.435 0.465 0.467 0.380 0.421 0.394 0.420 0.350 0.389 0.368 0.417 0.380 0.421
720 0.420 0.441 0.455 0.459 0.487 0.435 0.521 0.515 0.423 0.466 0.426 0.454 0.418 0.439 0.434 0.460 0.424 0.446
ETTh2
Avg 0.365 0.402 0.400 0.416 0.402 0.411 0.428 0.441 0.364 0.408 0.369 0.407 0.355 0.391 0.362 0.405 0.359 0.403
96 0.166 0.213 0.157 0.201 0.156 0.206 0.289 0.331 0.168 0.228 0.172 0.221 0.177 0.216 0.162 0.213 0.184 0.224
192 0.212 0.254 0.244 0.273 0.229 0.274 0.314 0.349 0.226 0.262 0.194 0.241 0.226 0.258 0.197 0.246 0.230 0.263
336 0.267 0.294 0.283 0.303 0.282 0.316 0.339 0.363 0.257 0.303 0.286 0.282 0.279 0.588 0.281 0.299 0.285 0.302
720 0.342 0.344 0.364 0.355 0.395 0.401 0.375 0.388 0.331 0.355 0.337 0.332 0.355 0.346 0.333 0.339 0.362 0.352
Weather
Avg 0.247 0.276 0.262 0.283 0.265 0.299 0.329 0.358 0.245 0.287 0.247 0.269 0.259 0.352 0.243 0.274 0.265 0.285
96 0.162 0.253 0.151 0.245 0.137 0.221 0.150 0.244 0.153 0.247 0.149 0.242 0.176 0.261 0.149 0.251 0.186 0.272
192 0.173 0.264 0.168 0.259 0.158 0.243 0.159 0.252 0.166 0.252 0.167 0.261 0.182 0.268 0.171 0.269 0.190 0.277
336 0.194 0.285 0.184 0.268 0.167 0.255 0.190 0.271 0.172 0.269 0.188 0.270 0.199 0.285 0.199 0.291 0.205 0.292
720 0.220 0.304 0.207 0.297 0.207 0.290 0.210 0.300 0.213 0.311 0.214 0.301 0.240 0.317 0.244 0.319 0.245 0.323
ECL
Avg 0.187 0.276 0.178 0.267 0.167 0.252 0.177 0.267 0.176 0.270 0.180 0.269 0.199 0.283 0.191 0.283 0.206 0.291
96 0.432 0.280 0.426 0.277 0.376 0.264 0.391 0.260 0.442 0.295 0.424 0.295 0.398 0.291 0.385 0.289 0.471 0.312
192 0.444 0.287 0.440 0.283 0.410 0.279 0.426 0.271 0.452 0.301 0.455 0.315 0.430 0.307 0.403 0.308 0.478 0.312
336 0.458 0.293 0.477 0.311 0.442 0.287 0.451 0.297 0.467 0.309 0.494 0.335 0.494 0.312 0.425 0.299 0.493 0.319
720 0.492 0.303 0.487 0.308 0.470 0.328 0.475 0.307 0.489 0.316 0.513 0.394 0.528 0.332 0.454 0.326 0.523 0.335
Traffic
Avg 0.457 0.291 0.458 0.295 0.424 0.289 0.436 0.284 0.463 0.305 0.471 0.334 0.463 0.310 0.417 0.306 0.491 0.320
96 0.084 0.201 0.083 0.199 0.089 0.211 0.098 0.228 0.091 0.214 0.090 0.209 0.082 0.200 0.147 0.279 0.087 0.218
192 0.174 0.295 0.190 0.307 0.175 0.289 0.196 0.325 0.185 0.307 0.188 0.310 0.172 0.295 0.234 0.354 0.171 0.294
336 0.331 0.416 0.401 0.458 0.345 0.423 0.359 0.433 0.345 0.414 0.342 0.427 0.329 0.415 0.403 0.474 0.349 0.418
720 0.847 0.694 0.879 0.702 0.882 0.744 0.875 0.713 0.874 0.729 0.885 0.707 0.889 0.747 1.103 0.804 0.873 0.713
Exchange
Avg 0.359 0.401 0.388 0.417 0.373 0.417 0.382 0.425 0.374 0.416 0.376 0.414 0.368 0.414 0.472 0.478 0.370 0.411
Average 0.336 0.349 0.355 0.358 0.346 0.361 0.372 0.378 0.339 0.357 0.341 0.357 0.344 0.364 0.348 0.366 0.345 0.359
Table 20: Full results for the long-term forecasting task compared with FilterNet FilterNet , TimesNet TimesNet , iTransformer iTransformer , PatchTST PatchTST , RLinear RLinear , DLinear DLinear and TimeMixer TimeMixer . (* means former, TNet is TimesNet, PTST is PatchTST, TMixer is TimeMixer) To ensure fairness in the comparison, we set the look-back window length of all models to 96. The standard deviation is within 0.5%. Red: best, Blue: second best.
{footnotesize }
SymTime FilterNet Chronos TNet iTrans* PTST RLinear DLinear TMixer
Methods (Ours) FilterNet Chronos TimesNet iTransformer PatchTST RLinear DLinear TimeMixer
Metrics MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
96 0.318 0.353 0.321 0.361 0.324 0.371 0.331 0.372 0.343 0.377 0.324 0.365 0.355 0.376 0.345 0.372 0.323 0.361
192 0.362 0.380 0.367 0.387 0.381 0.400 0.397 0.402 0.381 0.395 0.367 0.389 0.387 0.392 0.382 0.391 0.362 0.383
336 0.386 0.402 0.401 0.409 0.402 0.415 0.427 0.427 0.419 0.418 0.400 0.409 0.424 0.415 0.414 0.414 0.388 0.403
720 0.419 0.423 0.477 0.448 0.428 0.435 0.493 0.463 0.487 0.457 0.460 0.445 0.487 0.450 0.473 0.450 0.454 0.442
ETTm1
Avg 0.371 0.390 0.392 0.401 0.384 0.405 0.412 0.416 0.407 0.412 0.388 0.402 0.413 0.408 0.403 0.407 0.382 0.397
96 0.174 0.257 0.175 0.258 0.192 0.265 0.185 0.265 0.185 0.271 0.182 0.266 0.182 0.265 0.194 0.293 0.177 0.259
192 0.238 0.299 0.240 0.301 0.268 0.320 0.256 0.310 0.254 0.314 0.250 0.311 0.246 0.304 0.283 0.360 0.245 0.306
336 0.295 0.337 0.311 0.347 0.289 0.341 0.314 0.345 0.315 0.352 0.313 0.350 0.307 0.342 0.376 0.423 0.298 0.338
720 0.390 0.392 0.414 0.405 0.392 0.410 0.424 0.412 0.413 0.407 0.417 0.412 0.407 0.398 0.529 0.509 0.395 0.396
ETTm2
Avg 0.274 0.321 0.285 0.328 0.286 0.334 0.295 0.333 0.292 0.336 0.291 0.335 0.286 0.327 0.346 0.396 0.279 0.325
96 0.376 0.400 0.382 0.402 0.408 0.402 0.409 0.425 0.394 0.409 0.381 0.400 0.386 0.395 0.396 0.411 0.385 0.400
192 0.428 0.431 0.430 0.429 0.459 0.450 0.469 0.460 0.447 0.440 0.429 0.433 0.437 0.424 0.446 0.441 0.441 0.431
336 0.463 0.456 0.472 0.451 0.445 0.437 0.507 0.478 0.491 0.464 0.475 0.460 0.479 0.446 0.490 0.468 0.482 0.450
720 0.450 0.458 0.481 0.473 0.482 0.485 0.521 0.497 0.517 0.501 0.517 0.502 0.481 0.470 0.514 0.511 0.504 0.482
ETTh1
Avg 0.430 0.436 0.441 0.439 0.449 0.444 0.476 0.465 0.462 0.454 0.451 0.449 0.446 0.434 0.461 0.458 0.453 0.441
96 0.293 0.347 0.293 0.343 0.299 0.354 0.331 0.372 0.300 0.350 0.301 0.351 0.318 0.363 0.348 0.401 0.293 0.343
192 0.364 0.397 0.374 0.396 0.356 0.390 0.429 0.423 0.380 0.399 0.374 0.398 0.401 0.412 0.473 0.474 0.376 0.396
336 0.385 0.423 0.417 0.430 0.376 0.423 0.450 0.451 0.422 0.432 0.429 0.439 0.436 0.442 0.588 0.539 0.425 0.432
720 0.420 0.441 0.449 0.460 0.439 0.467 0.459 0.466 0.429 0.447 0.443 0.461 0.442 0.454 0.829 0.656 0.457 0.459
ETTh2
Avg 0.365 0.402 0.383 0.407 0.368 0.408 0.417 0.428 0.383 0.407 0.387 0.412 0.399 0.418 0.559 0.518 0.388 0.408
96 0.166 0.213 0.162 0.207 0.177 0.231 0.171 0.222 0.176 0.215 0.177 0.219 0.192 0.232 0.197 0.258 0.172 0.220
192 0.212 0.254 0.210 0.250 0.221 0.263 0.234 0.273 0.226 0.258 0.222 0.258 0.240 0.271 0.237 0.296 0.227 0.259
336 0.267 0.294 0.265 0.290 0.265 0.311 0.284 0.306 0.281 0.299 0.281 0.299 0.292 0.307 0.282 0.332 0.266 0.294
720 0.342 0.344 0.342 0.340 0.339 0.347 0.358 0.352 0.359 0.350 0.356 0.348 0.364 0.353 0.347 0.385 0.346 0.347
Weather
Avg 0.247 0.276 0.245 0.272 0.251 0.288 0.262 0.288 0.260 0.281 0.259 0.281 0.272 0.291 0.266 0.318 0.253 0.280
96 0.162 0.253 0.147 0.245 0.157 0.249 0.167 0.271 0.148 0.240 0.180 0.272 0.201 0.281 0.210 0.302 0.157 0.249
192 0.173 0.264 0.160 0.250 0.193 0.288 0.186 0.288 0.165 0.256 0.188 0.279 0.201 0.283 0.210 0.305 0.170 0.261
336 0.194 0.285 0.173 0.267 0.213 0.304 0.203 0.304 0.179 0.271 0.204 0.296 0.215 0.298 0.223 0.319 0.186 0.276
720 0.220 0.304 0.210 0.309 0.255 0.336 0.227 0.322 0.209 0.298 0.246 0.328 0.257 0.331 0.258 0.350 0.227 0.311
ECL
Avg 0.187 0.276 0.173 0.268 0.204 0.294 0.196 0.296 0.175 0.267 0.204 0.294 0.219 0.298 0.225 0.319 0.185 0.274
96 0.432 0.280 0.430 0.294 0.420 0.294 0.589 0.316 0.393 0.268 0.461 0.298 0.649 0.389 0.696 0.429 0.479 0.299
192 0.444 0.287 0.452 0.307 0.436 0.306 0.616 0.328 0.413 0.277 0.467 0.301 0.601 0.366 0.647 0.407 0.490 0.303
336 0.458 0.293 0.470 0.316 0.491 0.315 0.628 0.333 0.424 0.283 0.483 0.308 0.609 0.369 0.653 0.410 0.493 0.304
720 0.492 0.303 0.498 0.323 0.526 0.330 0.667 0.352 0.458 0.300 0.517 0.325 0.647 0.387 0.695 0.429 0.534 0.319
Traffic
Avg 0.457 0.291 0.463 0.310 0.468 0.312 0.625 0.332 0.422 0.282 0.482 0.308 0.627 0.378 0.673 0.419 0.499 0.306
96 0.084 0.201 0.091 0.211 0.090 0.207 0.115 0.246 0.094 0.216 0.088 0.205 0.093 0.217 0.093 0.226 0.091 0.210
192 0.174 0.295 0.186 0.305 0.190 0.316 0.213 0.335 0.185 0.307 0.189 0.309 0.184 0.307 0.184 0.324 0.185 0.304
336 0.331 0.416 0.380 0.449 0.354 0.419 0.367 0.440 0.336 0.422 0.327 0.415 0.351 0.432 0.328 0.436 0.361 0.435
720 0.847 0.694 0.896 0.712 0.892 0.716 0.978 0.753 0.893 0.716 0.886 0.706 0.886 0.714 0.880 0.705 0.974 0.741
Exchange
Avg 0.359 0.401 0.388 0.419 0.381 0.415 0.418 0.443 0.377 0.415 0.373 0.409 0.379 0.418 0.371 0.423 0.403 0.423
Average 0.336 0.349 0.346 0.356 0.349 0.363 0.388 0.375 0.347 0.357 0.354 0.361 0.380 0.371 0.413 0.407 0.355 0.357
Table 21: Full results for the long-term forecasting task compared with Autoformer Autoformer , Crossformer Crossformer , FEDformer FEDformer , ETSforemr ETSformer , Stationary Non-stationary-transformers , LightTS LightTS , Informer Informer . (Stationary means Nonstationary Transformer. * means former) To ensure fairness in the comparison, we set the look-back window length of all models to 96. The standard deviation is within 0.5%. Red: best, Blue: second best.
{footnotesize }
SymTime Autoformer Cross* FED* ETS* Stationary LightTS In*
Methods (Ours) Autoformer Crossformer FEDformer ETSformer Non-stationary-transformers LightTS Informer
Metrics MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
96 0.318 0.353 0.501 0.479 0.360 0.399 0.378 0.418 0.375 0.398 0.418 0.415 0.390 0.411 0.619 0.549
192 0.362 0.380 0.578 0.510 0.422 0.449 0.438 0.449 0.408 0.410 0.506 0.454 0.425 0.436 0.760 0.645
336 0.386 0.402 0.668 0.552 0.589 0.557 0.456 0.462 0.435 0.428 0.530 0.482 0.463 0.464 1.093 0.812
720 0.419 0.423 0.602 0.524 0.838 0.706 0.530 0.498 0.499 0.462 0.610 0.525 0.547 0.520 1.114 0.806
ETTm1
Avg 0.371 0.390 0.587 0.516 0.552 0.528 0.450 0.457 0.429 0.425 0.516 0.469 0.456 0.458 0.896 0.703
96 0.174 0.257 0.245 0.323 0.274 0.268 0.196 0.284 0.189 0.280 0.240 0.308 0.226 0.323 0.467 0.533
192 0.238 0.299 0.289 0.345 0.366 0.380 0.264 0.325 0.275 0.319 0.428 0.402 0.361 0.421 0.742 0.664
336 0.295 0.337 0.342 0.378 0.437 0.453 0.324 0.363 0.314 0.357 0.521 0.449 0.474 0.488 1.184 0.825
720 0.390 0.392 0.441 0.429 0.506 0.623 0.434 0.428 0.414 0.413 0.602 0.501 0.760 0.631 4.039 1.530
ETTm1
Avg 0.274 0.321 0.329 0.369 0.396 0.431 0.305 0.350 0.298 0.342 0.448 0.415 0.455 0.466 1.608 0.888
96 0.376 0.400 0.453 0.459 0.462 0.473 0.376 0.417 0.494 0.479 0.550 0.503 0.448 0.450 0.926 0.741
192 0.428 0.431 0.481 0.470 0.495 0.484 0.431 0.454 0.538 0.504 0.655 0.569 0.503 0.483 0.968 0.757
336 0.463 0.456 0.519 0.495 0.693 0.626 0.461 0.469 0.574 0.521 0.791 0.639 0.554 0.513 1.144 0.849
720 0.450 0.458 0.510 0.508 0.668 0.599 0.502 0.499 0.562 0.535 0.797 0.652 0.627 0.578 1.214 0.880
ETTh1
Avg 0.430 0.436 0.491 0.483 0.580 0.545 0.442 0.460 0.542 0.510 0.698 0.591 0.533 0.506 1.063 0.807
96 0.293 0.347 0.383 0.416 0.367 0.347 0.346 0.390 0.340 0.391 0.417 0.432 0.417 0.448 3.132 1.425
192 0.364 0.397 0.479 0.467 0.450 0.459 0.428 0.439 0.430 0.439 0.529 0.486 0.546 0.520 5.552 1.957
336 0.385 0.423 0.476 0.481 0.532 0.521 0.469 0.474 0.485 0.479 0.591 0.517 0.619 0.554 4.926 1.873
720 0.420 0.441 0.494 0.503 0.614 0.633 0.473 0.486 0.500 0.497 0.601 0.531 0.972 0.704 4.201 1.741
ETTh2
Avg 0.365 0.402 0.458 0.467 0.491 0.490 0.429 0.447 0.439 0.452 0.534 0.491 0.639 0.556 4.453 1.749
96 0.166 0.213 0.276 0.343 0.174 0.243 0.218 0.299 0.197 0.281 0.184 0.233 0.174 0.235 0.357 0.415
192 0.212 0.254 0.305 0.361 0.235 0.307 0.281 0.344 0.237 0.312 0.248 0.286 0.218 0.276 0.458 0.456
336 0.267 0.294 0.372 0.405 0.277 0.342 0.337 0.375 0.298 0.353 0.337 0.349 0.267 0.316 0.520 0.501
720 0.342 0.344 0.430 0.437 0.369 0.407 0.423 0.429 0.352 0.288 0.399 0.385 0.353 0.366 0.926 0.705
Weather
Avg 0.247 0.276 0.346 0.387 0.264 0.325 0.315 0.362 0.271 0.309 0.292 0.313 0.253 0.298 0.565 0.519
96 0.162 0.253 0.198 0.313 0.146 0.249 0.202 0.314 0.187 0.304 0.167 0.270 0.211 0.313 0.342 0.423
192 0.173 0.264 0.218 0.329 0.163 0.262 0.211 0.323 0.199 0.315 0.183 0.284 0.223 0.326 0.360 0.442
336 0.194 0.285 0.253 0.352 0.198 0.296 0.222 0.335 0.212 0.329 0.194 0.295 0.243 0.346 0.365 0.445
720 0.220 0.304 0.265 0.367 0.245 0.346 0.272 0.373 0.233 0.345 0.224 0.321 0.277 0.371 0.412 0.469
ECL
Avg 0.187 0.276 0.233 0.340 0.188 0.288 0.227 0.337 0.208 0.323 0.192 0.292 0.239 0.339 0.370 0.445
96 0.432 0.280 0.608 0.383 0.516 0.268 0.592 0.372 0.607 0.392 0.621 0.347 0.667 0.419 0.720 0.407
192 0.444 0.287 0.630 0.397 0.541 0.283 0.598 0.371 0.621 0.399 0.643 0.355 0.662 0.425 0.738 0.414
336 0.458 0.293 0.622 0.387 0.566 0.351 0.636 0.397 0.622 0.396 0.650 0.360 0.683 0.436 0.833 0.470
720 0.492 0.303 0.689 0.396 0.610 0.403 0.639 0.395 0.632 0.396 0.670 0.365 0.700 0.455 0.854 0.491
Traffic
Avg 0.457 0.291 0.637 0.391 0.558 0.326 0.616 0.384 0.621 0.396 0.646 0.357 0.678 0.434 0.786 0.445
96 0.084 0.201 0.191 0.318 0.276 0.383 0.162 0.291 0.085 0.204 0.132 0.254 0.128 0.266 0.896 0.761
192 0.174 0.295 0.315 0.407 0.540 0.552 0.276 0.382 0.182 0.303 0.251 0.361 0.292 0.402 1.146 0.861
336 0.331 0.416 0.480 0.519 1.229 0.873 0.442 0.488 0.348 0.428 0.467 0.507 0.500 0.536 1.628 1.017
720 0.847 0.694 1.255 0.868 1.721 1.055 1.175 0.833 1.025 0.774 1.304 0.837 1.002 0.763 2.552 1.299
Exchange
Avg 0.359 0.401 0.560 0.528 0.942 0.716 0.514 0.498 0.410 0.427 0.538 0.490 0.480 0.492 1.555 0.984
Average 0.336 0.349 0.455 0.435 0.496 0.456 0.456 0.412 0.402 0.398 0.483 0.427 0.467 0.444 1.412 0.818
Table 22: Full results for the long-term forecasting task compared with PatchTST PatchTST , TimeMixer TimeMixer , TimesNet TimesNet , Autoformer Autoformer , DLinear DLinear , iTransformer iTransformer , TimeXer TimeXer , FEDformer FEDformer . (* means former, PTST is PatchTST, TMixer is TimeMixer.) To ensure fairness in the comparison, we set the look-back window length of all models to 336. The standard deviation is within 0.5%. Red: best, Blue: second best.
{footnotesize }
Methods SymTime PTST TMixer TimesNet Auto* DLinear iTrans* TimeXer FED*
(Our) PatchTST TimeMixer TimesNet Autoformer DLinear iTransformer TimeXer FEDformer
Metrics MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
96 0.295 0.346 0.292 0.343 0.322 0.358 0.344 0.378 0.514 0.506 0.300 0.344 0.305 0.359 0.318 0.361 0.382 0.427
192 0.328 0.365 0.331 0.369 0.342 0.375 0.456 0.426 0.576 0.520 0.335 0.365 0.345 0.382 0.354 0.383 0.393 0.434
336 0.366 0.395 0.365 0.392 0.371 0.395 0.426 0.432 0.703 0.564 0.369 0.386 0.377 0.401 0.381 0.402 0.445 0.459
720 0.411 0.422 0.420 0.425 0.437 0.440 0.459 0.455 0.678 0.568 0.425 0.420 0.444 0.439 0.434 0.433 0.543 0.490
ETTm1
Avg 0.350 0.382 0.352 0.382 0.368 0.392 0.421 0.423 0.618 0.539 0.357 0.379 0.368 0.395 0.372 0.395 0.441 0.452
96 0.165 0.255 0.165 0.255 0.176 0.259 0.184 0.272 0.349 0.400 0.169 0.266 0.174 0.265 0.169 0.254 0.260 0.336
192 0.221 0.293 0.220 0.292 0.232 0.298 0.243 0.309 0.507 0.474 0.235 0.316 0.247 0.313 0.237 0.302 0.291 0.354
336 0.275 0.329 0.278 0.329 0.280 0.329 0.310 0.351 0.328 0.375 0.305 0.366 0.294 0.345 0.284 0.333 0.325 0.366
720 0.365 0.387 0.368 0.385 0.359 0.387 0.393 0.405 0.418 0.433 0.457 0.463 0.374 0.394 0.360 0.381 0.423 0.451
ETTm2
Avg 0.256 0.316 0.258 0.315 0.262 0.318 0.282 0.334 0.400 0.420 0.291 0.353 0.272 0.329 0.262 0.317 0.325 0.377
96 0.372 0.399 0.382 0.405 0.379 0.403 0.423 0.437 0.536 0.498 0.375 0.399 0.397 0.416 0.403 0.421 0.387 0.434
192 0.409 0.427 0.414 0.421 0.415 0.423 0.481 0.481 0.562 0.533 0.413 0.424 0.442 0.448 0.440 0.440 0.431 0.458
336 0.430 0.440 0.431 0.435 0.453 0.449 0.488 0.477 0.551 0.533 0.438 0.444 0.459 0.459 0.495 0.488 0.471 0.480
720 0.440 0.463 0.449 0.466 0.473 0.473 0.548 0.523 0.670 0.590 0.475 0.495 0.503 0.506 0.633 0.583 0.512 0.516
ETTh1
Avg 0.413 0.432 0.419 0.432 0.430 0.437 0.485 0.480 0.580 0.539 0.425 0.440 0.450 0.457 0.493 0.483 0.450 0.472
96 0.271 0.341 0.274 0.336 0.284 0.350 0.378 0.421 0.509 0.527 0.307 0.370 0.307 0.363 0.313 0.365 0.394 0.457
192 0.334 0.378 0.339 0.379 0.359 0.397 0.409 0.439 0.711 0.641 0.402 0.431 0.393 0.413 0.375 0.404 0.426 0.456
336 0.359 0.403 0.331 0.381 0.388 0.422 0.410 0.439 0.574 0.569 0.489 0.485 0.428 0.437 0.400 0.429 0.420 0.461
720 0.398 0.436 0.379 0.422 0.555 0.534 0.440 0.461 0.856 0.679 0.761 0.620 0.433 0.452 0.412 0.443 0.478 0.495
ETTh2
Avg 0.341 0.390 0.331 0.379 0.396 0.425 0.409 0.440 0.663 0.604 0.490 0.476 0.390 0.416 0.375 0.410 0.430 0.467
96 0.149 0.199 0.227 0.273 0.175 0.224 0.170 0.228 0.268 0.343 0.174 0.234 0.162 0.210 0.169 0.203 0.217 0.296
192 0.192 0.239 0.200 0.245 0.196 0.243 0.214 0.263 0.431 0.465 0.217 0.276 0.207 0.251 0.243 0.265 0.288 0.342
336 0.245 0.282 0.259 0.298 0.243 0.280 0.272 0.301 0.559 0.498 0.262 0.313 0.257 0.291 0.322 0.318 0.340 0.382
720 0.321 0.337 0.346 0.353 0.325 0.346 0.343 0.353 0.506 0.495 0.328 0.370 0.327 0.337 0.414 0.373 0.405 0.430
Weather
Avg 0.238 0.273 0.258 0.292 0.235 0.273 0.250 0.286 0.441 0.450 0.245 0.298 0.238 0.272 0.287 0.290 0.313 0.363
96 0.133 0.230 0.131 0.361 0.144 0.244 0.174 0.278 0.205 0.322 0.147 0.249 0.132 0.227 0.155 0.235 0.193 0.308
192 0.150 0.244 0.154 0.251 0.152 0.242 0.192 0.292 0.219 0.333 0.160 0.261 0.153 0.248 0.161 0.281 0.201 0.315
336 0.163 0.262 0.164 0.262 0.172 0.261 0.198 0.299 0.227 0.340 0.176 0.278 0.173 0.267 0.187 0.279 0.214 0.329
720 0.208 0.298 0.210 0.301 0.207 0.293 0.222 0.318 0.294 0.390 0.196 0.288 0.194 0.287 0.183 0.273 0.246 0.355
ECL
Avg 0.164 0.258 0.165 0.294 0.169 0.260 0.197 0.297 0.236 0.346 0.170 0.269 0.163 0.257 0.172 0.267 0.214 0.327
96 0.361 0.257 0.365 0.256 0.371 0.256 0.589 0.321 0.675 0.414 0.431 0.307 0.367 0.278 0.422 0.268 0.587 0.366
192 0.382 0.258 0.383 0.258 0.401 0.271 0.605 0.322 0.672 0.411 0.443 0.312 0.414 0.284 0.433 0.280 0.604 0.373
336 0.395 0.271 0.398 0.271 0.408 0.267 0.621 0.338 0.667 0.408 0.456 0.319 0.399 0.280 0.454 0.278 0.621 0.383
720 0.424 0.281 0.438 0.288 0.464 0.292 0.647 0.344 0.689 0.421 0.528 0.343 0.422 0.290 0.498 0.299 0.626 0.382
Traffic
Avg 0.391 0.267 0.396 0.268 0.411 0.271 0.615 0.331 0.676 0.413 0.465 0.320 0.401 0.283 0.452 0.281 0.610 0.376
96 0.085 0.204 0.093 0.214 0.092 0.217 0.218 0.343 0.967 0.778 0.099 0.235 0.099 0.224 0.234 0.312 0.168 0.285
192 0.180 0.301 0.200 0.321 0.235 0.345 0.299 0.411 0.931 0.750 0.195 0.335 0.215 0.338 0.283 0.402 0.186 0.296
336 0.335 0.419 0.373 0.448 0.370 0.441 0.468 0.527 1.051 0.809 0.380 0.474 0.378 0.454 0.433 0.343 0.250 0.342
720 0.869 0.700 0.875 0.695 0.963 0.750 1.208 0.847 1.261 0.893 1.120 0.805 0.876 0.690 0.688 0.942 0.899 0.784
Exchange
Avg 0.367 0.406 0.385 0.420 0.415 0.438 0.548 0.532 1.053 0.807 0.448 0.462 0.392 0.427 0.409 0.500 0.376 0.427
Average 0.315 0.341 0.320 0.348 0.336 0.352 0.401 0.390 0.583 0.515 0.361 0.375 0.347 0.355 0.353 0.368 0.395 0.408
Table 23: Full results for the long-term forecasting task compared with PatchTST PatchTST , TimeMixer TimeMixer , TimesNet TimesNet , Autoformer Autoformer , DLinear DLinear , iTransformer iTransformer , TimeXer TimeXer , FITS FITS . (* means former, TMixer is TimeMixer.) To ensure fairness in the comparison, we set the look-back window length of all models to 512. The standard deviation is within 0.5%. Red: best, Blue: second best.
{footnotesize }
SymTime PatchTST TMixer TimesNet Auto* DLinear iTrans* TimeXer FITS
Methods (Ours) PatchTST TimeMixer TimesNet Autoformer DLinear iTransformer TimeXer FITS
Metrics MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
96 0.313 0.348 0.290 0.344 0.315 0.354 0.358 0.388 0.495 0.499 0.304 0.347 0.310 0.364 0.332 0.374 0.306 0.349
192 0.326 0.363 0.333 0.371 0.344 0.376 0.457 0.439 0.549 0.514 0.337 0.368 0.349 0.387 0.364 0.392 0.338 0.367
336 0.376 0.390 0.369 0.392 0.385 0.397 0.412 0.427 0.536 0.509 0.366 0.385 0.376 0.401 0.387 0.407 0.421 0.384
720 0.409 0.420 0.416 0.420 0.441 0.440 0.473 0.467 0.645 0.550 0.424 0.422 0.434 0.436 0.430 0.432 0.432 0.437
ETTm1
Avg 0.356 0.380 0.352 0.382 0.371 0.392 0.425 0.430 0.556 0.518 0.358 0.380 0.367 0.397 0.378 0.401 0.374 0.384
96 0.169 0.258 0.166 0.256 0.173 0.261 0.192 0.279 0.278 0.353 0.166 0.262 0.182 0.272 0.177 0.263 0.165 0.254
192 0.232 0.301 0.223 0.296 0.223 0.298 0.264 0.325 0.318 0.383 0.225 0.304 0.242 0.312 0.241 0.312 0.219 0.291
336 0.287 0.333 0.274 0.329 0.291 0.342 0.320 0.364 0.411 0.444 0.299 0.361 0.292 0.346 0.294 0.340 0.272 0.326
720 0.372 0.387 0.362 0.385 0.366 0.391 0.402 0.410 0.476 0.484 0.412 0.432 0.378 0.396 0.382 0.399 0.359 0.391
ETTm2
Avg 0.265 0.320 0.256 0.317 0.263 0.323 0.294 0.344 0.371 0.416 0.275 0.340 0.273 0.331 0.274 0.329 0.254 0.313
96 0.372 0.389 0.370 0.400 0.380 0.408 0.442 0.465 0.556 0.534 0.368 0.397 0.394 0.420 0.397 0.424 0.372 0.396
192 0.402 0.424 0.413 0.429 0.431 0.444 0.473 0.475 0.568 0.538 0.400 0.417 0.430 0.444 0.438 0.452 0.405 0.415
336 0.433 0.445 0.422 0.440 0.468 0.472 0.505 0.501 0.606 0.573 0.430 0.442 0.447 0.459 0.466 0.472 0.440 0.468
720 0.448 0.469 0.447 0.468 0.435 0.453 0.505 0.504 0.777 0.669 0.476 0.497 0.514 0.516 0.597 0.566 0.453 0.485
ETTh1
Avg 0.414 0.432 0.413 0.434 0.429 0.444 0.481 0.486 0.627 0.579 0.418 0.438 0.446 0.460 0.475 0.479 0.418 0.441
96 0.290 0.352 0.273 0.337 0.298 0.364 0.336 0.391 0.447 0.486 0.288 0.355 0.317 0.367 0.299 0.358 0.291 0.353
192 0.376 0.409 0.341 0.382 0.361 0.399 0.393 0.425 0.587 0.568 0.394 0.427 0.388 0.411 0.370 0.408 0.350 0.395
336 0.382 0.416 0.398 0.458 0.393 0.421 0.406 0.445 0.689 0.608 0.501 0.491 0.422 0.437 0.372 0.410 0.375 0.424
720 0.414 0.442 0.416 0.458 0.442 0.458 0.451 0.466 1.027 0.772 0.813 0.638 0.424 0.454 0.376 0.424 0.437 0.459
ETTh2
Avg 0.365 0.405 0.357 0.409 0.373 0.410 0.397 0.432 0.687 0.609 0.499 0.478 0.388 0.417 0.354 0.400 0.363 0.408
96 0.159 0.205 0.230 0.282 0.169 0.225 0.168 0.224 0.375 0.426 0.171 0.230 0.175 0.223 0.175 0.209 0.172 0.226
192 0.203 0.260 0.194 0.242 0.191 0.242 0.217 0.266 0.471 0.491 0.213 0.269 0.213 0.256 0.246 0.267 0.216 0.262
336 0.256 0.289 0.245 0.282 0.246 0.283 0.278 0.310 0.514 0.521 0.260 0.312 0.265 0.296 0.314 0.309 0.261 0.295
720 0.319 0.339 0.312 0.332 0.316 0.333 0.342 0.352 0.596 0.506 0.320 0.358 0.342 0.347 0.391 0.353 0.326 0.342
Weather
Avg 0.234 0.273 0.245 0.284 0.231 0.271 0.251 0.288 0.489 0.486 0.241 0.292 0.249 0.280 0.282 0.284 0.244 0.281
96 0.128 0.256 0.129 0.224 0.137 0.229 0.184 0.288 0.215 0.328 0.141 0.241 0.131 0.227 0.140 0.242 0.145 0.244
192 0.149 0.244 0.158 0.258 0.152 0.250 0.187 0.291 0.224 0.333 0.154 0.254 0.155 0.250 0.157 0.256 0.153 0.250
336 0.161 0.261 0.163 0.261 0.193 0.295 0.206 0.307 0.235 0.341 0.169 0.271 0.171 0.266 0.176 0.275 0.169 0.266
720 0.216 0.308 0.225 0.331 0.228 0.324 0.226 0.323 0.738 0.570 0.204 0.304 0.191 0.285 0.211 0.306 0.208 0.298
ECL
Avg 0.163 0.267 0.169 0.269 0.177 0.274 0.201 0.302 0.353 0.393 0.167 0.267 0.162 0.257 0.171 0.270 0.169 0.265
96 0.365 0.257 0.369 0.262 0.368 0.254 0.600 0.321 0.697 0.428 0.412 0.294 0.351 0.257 0.428 0.271 0.398 0.277
192 0.379 0.256 0.379 0.253 0.399 0.268 0.612 0.328 0.700 0.429 0.422 0.299 0.373 0.268 0.448 0.282 0.409 0.280
336 0.401 0.275 0.410 0.280 0.404 0.264 0.631 0.338 0.707 0.437 0.431 0.304 0.386 0.274 0.473 0.289 0.418 0.285
720 0.433 0.284 0.438 0.291 0.467 0.293 0.654 0.352 0.718 0.445 0.468 0.325 0.424 0.294 0.516 0.307 0.456 0.306
Traffic
Avg 0.395 0.268 0.399 0.272 0.410 0.270 0.624 0.334 0.705 0.435 0.433 0.305 0.383 0.273 0.466 0.287 0.420 0.287
96 0.089 0.212 0.095 0.220 0.096 0.220 0.238 0.366 0.617 0.623 0.120 0.260 0.135 0.267 0.099 0.223 0.100 0.225
192 0.176 0.298 0.215 0.336 0.197 0.318 0.439 0.497 0.810 0.739 0.241 0.376 0.322 0.417 0.207 0.331 0.201 0.326
336 0.380 0.439 0.392 0.459 0.737 0.642 0.647 0.620 0.858 0.746 0.439 0.509 0.361 0.448 0.767 0.689 0.350 0.437
720 0.893 0.698 0.890 0.680 1.038 0.808 1.550 0.948 1.491 0.965 1.199 0.831 0.888 0.736 0.984 0.780 0.920 0.728
Exchange
Avg 0.384 0.412 0.398 0.423 0.517 0.497 0.718 0.608 0.944 0.768 0.500 0.494 0.427 0.467 0.514 0.506 0.393 0.439
Average 0.322 0.345 0.324 0.349 0.346 0.360 0.424 0.403 0.591 0.525 0.361 0.374 0.337 0.360 0.364 0.369 0.329 0.352
Table 24: Full results for the short-term forecasting task in the M4 dataset compared with Peri-midFormer Peri-midFormer , $S^{2}$IP-LLM(S-LLM) S2IP-LLM , Time-LLM(T-LLM) Time-LLM , GPT4TS GPT4TS , TimeMixer TimeMixer , PatchTST PatchTST , iTransformer iTransformer , TimesNet TimesNet , DLinear DLinear , Informer Informer . (* means former.) The standard deviation is within 0.5%. Red: best, Blue: second best.
{footnotesize }
Methods SymTime Peri-mid* S-LLM T-LLM GPT4TS TimeMixer PatchTST iTrans* TimesNet DLinear In*
Metric (Ours) Peri-midFormer S2IP-LLM Time-LLM GPT4TS TimeMixer PatchTST iTransformer TimesNet DLinear Informer
SMAPE 13.355 13.483 14.931 13.450 14.847 13.369 13.677 13.724 13.463 14.340 14.698
MASE 2.997 3.080 3.345 3.184 3.628 3.009 3.049 3.157 3.058 3.112 3.293
Yearly
OWA 0.786 0.800 0.878 0.819 0.911 0.787 0.802 0.817 0.797 0.830 0.864
SMAPE 10.060 10.037 10.655 10.671 10.389 10.131 10.922 13.473 10.069 10.510 16.172
MASE 1.183 1.170 1.249 1.276 1.228 1.186 1.326 1.722 1.175 1.241 2.136
Quarterly
OWA 0.872 0.882 0.939 0.950 0.919 0.893 0.979 1.240 0.886 0.930 1.513
SMAPE 12.608 12.795 13.012 13.416 12.907 12.762 14.200 13.674 12.760 13.382 15.446
MASE 0.925 0.948 0.973 1.045 0.954 0.940 1.111 1.068 0.947 1.007 1.247
Monthly
OWA 0.872 0.889 0.909 0.957 0.896 0.884 1.015 0.976 0.887 0.937 1.122
SMAPE 4.941 4.912 5.540 4.973 5.266 5.085 5.658 5.598 4.995 5.122 6.839
MASE 3.327 3.260 8.426 3.412 3.595 3.403 3.626 3.957 3.346 3.608 4.536
Others
OWA 1.045 1.031 3.792 1.059 1.121 1.072 1.167 1.213 1.053 1.108 1.435
SMAPE 11.785 11.897 12.514 12.584 12.367 11.885 12.866 13.233 11.888 12.500 15.018
MASE 1.584 1.607 1.726 1.763 1.767 1.598 1.734 1.850 1.607 1.678 2.096
Average
OWA 0.849 0.859 0.913 0.915 0.918 0.856 0.928 0.972 0.858 0.899 1.102
Table 25: Full results for the short-term forecasting task in the M4 dataset compared with LightTS LightTS , Autoformer Autoformer , Crossformer Crossformer , FEDformer FEDformer , ETSformer ETSformer , Nonstationary Transformer (Stationary) Non-stationary-transformers , FiLM FiLM , MICN MICN , Reformer Reformer , Pyraformer Pyraformer . The standard deviation is within 0.5%. (* means former.) Red: best, Blue: second best.
{footnotesize }
Methods SymTime LightTS Auto* Cross* FED* ETS* Stationary FiLM MICN Re* Pyra*
Metric (Ours) LightTS Autoformer Crossformer FEDformer ETSformer Non-stationary-transformers FiLM MICN Reformer Pyraformer
SMAPE 13.355 13.444 17.764 79.308 13.508 18.009 13.717 14.076 14.557 13.752 14.594
MASE 2.997 3.022 3.919 18.692 3.051 4.487 3.078 3.017 3.380 3.088 3.269
Yearly
OWA 0.786 0.792 1.037 4.778 0.797 1.115 0.807 0.810 0.871 0.809 0.858
SMAPE 10.060 10.252 13.968 74.943 10.706 13.376 10.958 10.711 11.408 10.900 11.654
MASE 1.183 1.183 1.754 13.133 1.263 1.906 1.325 1.292 1.384 1.316 1.392
Quarterly
OWA 0.872 0.897 1.274 8.191 0.947 1.302 0.981 0.957 1.022 0.975 1.037
SMAPE 12.608 12.798 18.200 68.892 13.925 14.588 13.917 13.362 13.803 13.949 14.963
MASE 0.925 0.957 1.574 11.199 1.062 1.368 1.097 1.016 1.078 1.096 1.165
Monthly
OWA 0.872 0.894 1.371 7.654 0.982 1.149 0.998 0.941 0.985 0.999 1.066
SMAPE 4.941 5.324 6.738 176.164 4.888 7.267 6.302 5.387 6.090 6.611 5.605
MASE 3.327 3.410 4.853 116.723 3.244 5.240 4.064 3.670 4.203 4.492 3.966
Others
OWA 1.045 1.098 1.474 36.941 1.026 1.591 1.304 1.146 1.304 1.404 1.215
SMAPE 11.785 11.962 16.511 78.103 12.605 14.718 12.780 12.491 13.016 12.805 13.616
MASE 1.584 1.609 2.321 18.663 1.677 2.408 1.756 1.675 1.837 1.777 1.843
Average
OWA 0.849 0.862 1.215 7.759 0.903 1.172 0.930 0.899 0.960 0.937 0.984
Table 26: Full results for time series classification task compared with (1) classical methods: DTW DTW , XGBoost XGBoost , Rocket Rocket ; (2) RNN-based methods: LSTM LSTM , LSTNet LSTNet , LSSL LSSL ; (3) CNN-based methods: InceptionTime (InTime) InceptionTime , TCN TCN , TimesNet TimesNet , TSLANet TSLANet . We report the classification accuracy (%) as the result. Red: best, Blue: second best. The standard deviation is within 1%.
{footnotesize }
Classical Methods RNN-based CNN-based
DTW XGBoost Rocket LSTM LSTNet LSSL InTime TCN TimesNet TSLANet SymTime
Datasets / Methods DTW XGBoost Rocket LSTM LSTNet LSSL InceptionTime TCN TimesNet TSLANet (Ours)
EthanolConcentration 32.3 43.7 45.2 32.3 39.9 31.1 39.1 28.9 35.7 30.4 37.3
FaceDetection 52.9 63.3 64.7 57.7 65.7 66.7 65.4 52.8 68.6 66.7 69.2
Handwriting 28.6 15.8 58.8 15.2 25.8 24.6 46.9 53.3 32.1 57.9 36.7
Heartbeat 71.7 73.2 75.6 72.2 77.1 72.7 75.2 75.6 78.0 77.5 74.1
JapaneseVowels 94.9 86.5 96.2 79.7 98.1 98.4 95.1 98.9 98.4 95.1 98.1
PEMS-SF 71.1 98.3 75.1 39.9 86.7 86.1 79.6 68.8 89.6 83.8 97.1
SelfRegulationSCP1 77.7 84.6 90.8 68.9 84.0 90.8 87.2 84.6 91.8 91.8 89.8
SelfRegulationSCP2 53.9 48.9 53.3 46.6 52.8 52.2 53.6 55.6 57.2 53.3 58.9
SpokenArabicDigits 96.3 69.6 71.2 31.9 100.0 100.0 96.3 95.6 99.0 98.0 98.9
UWaveGestureLibrary 90.3 75.9 94.4 41.2 87.8 85.9 92.4 88.4 85.3 89.4 89.4
Average Accuracy 67.0 66.0 72.5 48.6 71.8 70.9 73.1 70.3 73.6 74.4 74.9
Table 27: Full reuslts for time series classification task compared with (1) Transformer-based methods: Autoformer Autoformer , FEDformer FEDformer , ETSformer ETSformer , Informer Informer , iTransformer iTransformer , PatchTST (Patch) PatchTST , GPT4TS (GPT) GPT4TS , UniTS UniTS , Peri-midformer Peri-midFormer and (2) MLP-based methods: DLinear DLinear , LightTS LightTS . We report the classification accuracy (%) as the results. (* means former.) Red: best, Blue: second best. The standard deviation is within 1%.
{footnotesize }
Transformer-based MLP-based
Auto* FED* ETS* In* iTrans* Patch GPT UniTS Peri-mid* DLinear LightTS SymTime
Datasets / Methods Autoformer FEDformer ETSformer Informer iTransformer PatchTST GPT4TS UniTS Peri-midFormer DLinear LightTS (Ours)
EthanolConcentration 31.6 31.2 28.1 31.6 27.0 29.6 34.2 37.3 47.3 32.6 29.7 37.3
FaceDetection 68.4 66.0 66.3 67.0 67.0 67.8 69.2 67.5 68.7 68.0 67.5 69.2
Handwriting 36.7 28.0 32.5 32.8 27.2 23.2 32.7 27.0 31.5 27.0 26.1 36.7
Heartbeat 74.6 73.7 71.2 80.5 75.6 75.7 77.2 80.5 86.3 75.1 75.1 74.1
JapaneseVowels 96.2 98.4 95.9 98.9 97.6 94.0 98.6 97.8 96.8 96.2 96.2 98.1
PEMS-SF 82.7 80.9 86.0 81.5 85.5 80.9 87.9 93.1 88.2 75.1 88.4 97.1
SelfRegulationSCP1 84.0 88.7 89.6 90.1 92.2 82.2 87.2 89.6 87.4 87.3 89.8 89.8
SelfRegulationSCP2 50.6 54.4 55.0 53.3 54.4 53.6 59.4 61.1 55.4 50.5 51.1 58.9
SpokenArabicDigits 100.0 100.0 100.0 100.0 98.0 98.0 95.2 98.9 98.0 81.4 100.0 98.9
UWaveGestureLibrary 85.9 85.3 85.0 85.6 85.9 81.7 85.1 87.8 84.3 82.1 80.3 89.4
Table 28: Full results for time series imputation task, where we randomly mask {12.5%, 25%, 37.5%, 50%} time points of length-96 time series to compare the model performance under different missing degrees. We compare with GPT4TS GPT4TS , TimesNet TimesNet , Peri-midFormer Peri-midFormer , Moment Moment , iTransformer iTransformer , PatchTST PatchTST , DLinear DLinear in this table. (* means former.) The standard deviation is within 0.5%. Red: best, Blue: second best.
{footnotesize }
SymTime GPT4TS TimesNet Peri-mid* Moment iTrans* PatchTST DLinear
Models (Ours) GPT4TS TimesNet Peri-midFormer Moment iTransformer PatchTST DLinear
Mask Ratio MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
12.5% 0.032 0.110 0.018 0.090 0.019 0.091 0.032 0.109 0.069 0.170 0.046 0.147 0.045 0.137 0.056 0.162
25% 0.034 0.113 0.024 0.102 0.024 0.101 0.034 0.112 0.071 0.169 0.060 0.171 0.046 0.139 0.077 0.191
37.5% 0.037 0.118 0.029 0.111 0.029 0.112 0.037 0.117 0.069 0.163 0.077 0.195 0.049 0.143 0.100 0.218
50% 0.041 0.126 0.042 0.132 0.036 0.124 0.042 0.126 0.086 0.169 0.104 0.228 0.055 0.152 0.129 0.247
ETTm1
Avg 0.036 0.116 0.028 0.109 0.027 0.107 0.036 0.116 0.074 0.168 0.072 0.185 0.049 0.143 0.090 0.204
12.5% 0.024 0.084 0.018 0.081 0.019 0.081 0.023 0.081 0.032 0.108 0.052 0.151 0.026 0.094 0.067 0.171
25% 0.024 0.086 0.021 0.082 0.021 0.086 0.024 0.084 0.029 0.105 0.070 0.179 0.028 0.099 0.089 0.200
37.5% 0.027 0.089 0.023 0.090 0.023 0.091 0.026 0.089 0.032 0.109 0.091 0.204 0.031 0.104 0.112 0.226
50% 0.030 0.093 0.027 0.098 0.026 0.098 0.030 0.095 0.031 0.110 0.117 0.232 0.034 0.109 0.140 0.253
ETTm2
Avg 0.026 0.088 0.022 0.088 0.022 0.089 0.026 0.087 0.031 0.108 0.082 0.191 0.030 0.101 0.102 0.212
12.5% 0.074 0.179 0.063 0.171 0.062 0.169 0.069 0.173 0.160 0.239 0.098 0.220 0.097 0.203 0.111 0.232
25% 0.082 0.190 0.080 0.190 0.081 0.191 0.079 0.185 0.142 0.238 0.125 0.249 0.115 0.221 0.149 0.269
37.5% 0.100 0.205 0.107 0.218 0.098 0.210 0.096 0.202 0.121 0.228 0.156 0.278 0.134 0.239 0.187 0.301
50% 0.123 0.230 0.121 0.221 0.116 0.227 0.122 0.226 0.132 0.231 0.213 0.327 0.160 0.260 0.229 0.332
ETTh1
Avg 0.095 0.201 0.093 0.200 0.089 0.199 0.091 0.196 0.139 0.234 0.148 0.269 0.126 0.231 0.169 0.283
12.5% 0.051 0.138 0.041 0.129 0.040 0.132 0.051 0.139 0.051 0.150 0.095 0.210 0.058 0.153 0.109 0.223
25% 0.055 0.146 0.046 0.138 0.047 0.144 0.054 0.142 0.079 0.177 0.120 0.239 0.063 0.160 0.146 0.260
37.5% 0.059 0.152 0.060 0.160 0.054 0.154 0.058 0.148 0.056 0.155 0.149 0.266 0.068 0.167 0.180 0.290
50% 0.064 0.157 0.061 0.160 0.061 0.164 0.064 0.159 0.056 0.154 0.192 0.302 0.074 0.175 0.217 0.319
ETTh2
Avg 0.058 0.148 0.052 0.147 0.050 0.148 0.057 0.147 0.061 0.159 0.139 0.254 0.066 0.164 0.163 0.273
12.5% 0.037 0.122 0.080 0.195 0.088 0.203 0.047 0.140 0.095 0.211 0.073 0.190 0.061 0.170 0.084 0.206
25% 0.046 0.139 0.089 0.205 0.092 0.208 0.053 0.162 0.093 0.211 0.090 0.214 0.072 0.185 0.113 0.243
37.5% 0.060 0.160 0.094 0.217 0.096 0.214 0.067 0.179 0.094 0.211 0.107 0.235 0.082 0.198 0.141 0.273
50% 0.075 0.181 0.108 0.231 0.102 0.221 0.085 0.195 0.092 0.210 0.127 0.257 0.097 0.216 0.173 0.303
ECL
Avg 0.054 0.151 0.093 0.212 0.094 0.211 0.063 0.169 0.094 0.211 0.099 0.224 0.078 0.192 0.128 0.256
12.5% 0.025 0.035 0.026 0.047 0.026 0.049 0.025 0.037 0.033 0.073 0.038 0.087 0.028 0.049 0.039 0.091
25% 0.027 0.037 0.030 0.055 0.030 0.056 0.026 0.037 0.036 0.078 0.046 0.106 0.032 0.055 0.049 0.112
37.5% 0.029 0.039 0.033 0.061 0.032 0.058 0.029 0.041 0.034 0.075 0.055 0.122 0.035 0.059 0.057 0.125
50% 0.032 0.042 0.039 0.070 0.034 0.062 0.034 0.048 0.035 0.075 0.068 0.142 0.039 0.064 0.067 0.139
Weather
Avg 0.028 0.038 0.032 0.058 0.030 0.056 0.029 0.041 0.035 0.075 0.052 0.114 0.033 0.057 0.053 0.116
Average 0.049 0.124 0.053 0.136 0.052 0.135 0.050 0.126 0.072 0.159 0.099 0.206 0.064 0.148 0.118 0.224
Table 29: Full results for time series imputation task, where we randomly mask {12.5%, 25%, 37.5%, 50%} time points of length-96 time series to compare the model performance under different missing degrees. We compare with Stationary Non-stationary-transformers , LightTS LightTS , ETSformer ETSformer , FEDformer FEDformer , Informer Informer , Reformer Reformer and Pyraformer Pyraformer in this table. (Stationary means Nonstationary Transformer.) The standard deviation is within 0.5%. Red: best, Blue: second best.
{footnotesize }
SymTime Stationary LightTS ETSformer FEDformer Informer Reformer Pyraformer
Methods (Ours) Non-stationary-transformers LightTS ETSformer FEDformer Informer Reformer Pyraformer
Mask Ratio MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
12.5% 0.032 0.110 0.026 0.107 0.054 0.158 0.034 0.130 0.068 0.188 0.027 0.115 0.032 0.126 0.670 0.541
25% 0.034 0.113 0.032 0.119 0.061 0.173 0.053 0.162 0.097 0.230 0.040 0.140 0.042 0.146 0.689 0.553
37.5% 0.037 0.118 0.039 0.131 0.073 0.189 0.082 0.201 0.134 0.287 0.071 0.189 0.063 0.182 0.737 0.581
50% 0.041 0.126 0.047 0.145 0.086 0.207 0.130 0.257 0.188 0.323 0.091 0.208 0.082 0.208 0.770 0.605
ETTm1
Avg 0.036 0.116 0.036 0.126 0.068 0.182 0.075 0.187 0.121 0.257 0.057 0.163 0.055 0.166 0.717 0.570
12.5% 0.024 0.084 0.021 0.088 0.051 0.150 0.061 0.169 0.109 0.239 0.196 0.326 0.108 0.228 0.394 0.470
25% 0.024 0.086 0.024 0.096 0.069 0.176 0.093 0.214 0.166 0.295 0.295 0.414 0.136 0.262 0.421 0.482
37.5% 0.027 0.089 0.027 0.103 0.074 0.185 0.137 0.253 0.237 0.356 0.155 0.293 0.175 0.300 0.478 0.521
50% 0.030 0.093 0.030 0.108 0.078 0.192 0.237 0.332 0.323 0.412 0.214 0.325 0.211 0.329 0.568 0.560
ETTm2
Avg 0.026 0.088 0.026 0.099 0.068 0.176 0.132 0.242 0.209 0.326 0.215 0.340 0.157 0.280 0.465 0.508
12.5% 0.074 0.179 0.060 0.165 0.119 0.239 0.073 0.195 0.126 0.265 0.068 0.187 0.074 0.194 0.857 0.609
25% 0.082 0.190 0.080 0.189 0.144 0.266 0.105 0.234 0.169 0.305 0.096 0.220 0.102 0.227 0.829 0.672
37.5% 0.100 0.205 0.102 0.212 0.171 0.292 0.144 0.276 0.220 0.348 0.128 0.253 0.135 0.261 0.830 0.675
50% 0.123 0.230 0.133 0.240 0.201 0.317 0.200 0.327 0.298 0.403 0.166 0.287 0.179 0.298 0.854 0.691
ETTh1
Avg 0.095 0.201 0.094 0.201 0.159 0.278 0.130 0.258 0.204 0.330 0.115 0.237 0.122 0.245 0.842 0.682
12.5% 0.051 0.138 0.042 0.133 0.094 0.208 0.134 0.251 0.187 0.319 0.271 0.384 0.163 0.289 0.976 0.754
25% 0.055 0.146 0.049 0.147 0.140 0.255 0.180 0.294 0.279 0.396 0.362 0.450 0.206 0.331 1.037 0.774
37.5% 0.059 0.152 0.056 0.158 0.159 0.274 0.243 0.341 0.402 0.465 0.401 0.469 0.252 0.370 1.107 0.800
50% 0.064 0.157 0.065 0.170 0.180 0.293 0.353 0.408 0.604 0.504 0.437 0.487 0.316 0.419 1.193 0.838
ETTh2
Avg 0.058 0.148 0.053 0.152 0.143 0.258 0.228 0.324 0.368 0.421 0.368 0.448 0.234 0.352 1.079 0.792
12.5% 0.037 0.122 0.093 0.210 0.077 0.198 0.185 0.323 0.197 0.324 0.152 0.279 0.190 0.308 0.297 0.383
25% 0.046 0.139 0.097 0.214 0.099 0.228 0.207 0.340 0.208 0.345 0.166 0.290 0.197 0.312 0.294 0.380
37.5% 0.060 0.160 0.102 0.220 0.120 0.252 0.226 0.355 0.219 0.337 0.178 0.297 0.203 0.315 0.296 0.381
50% 0.075 0.181 0.108 0.228 0.138 0.272 0.251 0.372 0.235 0.357 0.189 0.305 0.210 0.319 0.299 0.383
ECL
Avg 0.049 0.151 0.100 0.218 0.108 0.238 0.217 0.347 0.215 0.341 0.171 0.293 0.200 0.313 0.297 0.382
12.5% 0.025 0.035 0.027 0.051 0.039 0.092 0.042 0.103 0.057 0.141 0.040 0.108 0.031 0.076 0.140 0.220
25% 0.027 0.037 0.029 0.056 0.045 0.105 0.056 0.131 0.066 0.155 0.045 0.130 0.035 0.082 0.147 0.229
37.5% 0.029 0.039 0.033 0.062 0.049 0.110 0.081 0.180 0.083 0.180 0.049 0.101 0.040 0.091 0.156 0.240
50% 0.032 0.042 0.037 0.068 0.054 0.117 0.102 0.207 0.103 0.207 0.054 0.114 0.046 0.099 0.164 0.249
Weather
Avg 0.028 0.038 0.032 0.059 0.047 0.106 0.071 0.155 0.077 0.171 0.047 0.113 0.038 0.087 0.152 0.235
Average 0.049 0.124 0.057 0.143 0.099 0.206 0.142 0.252 0.199 0.308 0.162 0.265 0.134 0.241 0.592 0.528
Table 30: Full reuslts for time series anomaly detection task, where P, R and F1 represent the precision, recall and F1-score (%) respectively. F1-score is the harmonic mean of precision and recall. A higher value of P, R and F1 indicates a better performance. We compare with: Transformer Transformer , Reformer Reformer , Informer Informer , Autoformer Autoformer , Crossformer Crossformer , iTransformer iTransformer , Anomaly Anomaly-Transformer , Stationary Non-stationary-transformers , DLinear DLinear , LightTS LightTS , ETSformer ETSformer , FEDformer FEDformer , PatchTST PatchTST , TimesNet TimesNet , GPT4TS GPT4TS , Peri-midFormer Peri-midFormer , UniTS UniTS , where Anomaly means the Anomaly Transformer and Stationary means the Non-stationary Transformer. The standard deviation is within 1%. Red: best, Blue: second best.
{footnotesize }
Datasets SMD MSL SMAP SWaT PSM Avg F1
Metircs P R F1 P R F1 P R F1 P R F1 P R F1 (%)
LSTM LSTM 78.52 65.47 71.41 78.04 86.22 81.93 91.06 57.49 70.48 78.06 91.72 84.34 69.24 99.53 81.67 77.97
LogTrans Transformer 83.46 70.13 76.21 73.05 87.37 79.57 89.15 57.59 69.97 68.67 97.32 80.52 63.06 98.00 76.74 76.60
Transformer Transformer 78.44 65.26 71.24 89.85 73.71 80.99 90.77 61.76 73.50 96.82 66.41 79.76 99.31 83.18 90.53 79.20
TCN TCN 84.06 79.07 81.49 75.11 82.44 78.60 86.90 59.23 70.45 76.59 95.71 85.09 54.59 99.77 70.57 77.24
LSSL LSSL 78.51 65.32 71.31 77.55 88.18 82.53 89.43 53.43 66.90 79.05 93.72 85.76 66.02 92.93 77.20 76.74
Reformer Reformer 72.50 84.19 77.90 90.24 73.78 81.18 90.63 62.48 73.97 99.94 66.75 80.04 99.73 83.03 90.62 80.74
Informer Informer 72.51 84.13 77.88 90.10 73.68 81.07 90.57 61.51 73.26 99.83 67.24 80.35 99.03 83.21 90.43 80.60
Autoformer Autoformer 78.46 65.11 71.17 90.59 75.26 82.22 90.84 62.39 73.97 99.95 65.57 79.19 99.99 78.96 88.24 78.96
Crossformer Crossformer 71.89 83.41 77.22 90.32 72.74 80.59 89.68 53.63 67.12 98.00 83.59 90.22 97.49 88.02 92.52 81.53
iTransformer iTransformer 76.13 84.70 80.19 86.15 62.54 72.47 90.68 52.78 66.72 92.23 93.05 92.64 97.92 92.03 94.88 81.38
Pyraformer Pyraformer 85.61 80.61 83.04 83.81 85.93 84.86 92.54 57.71 71.09 87.92 96.00 91.78 71.67 96.02 82.08 82.57
Anomaly Anomaly-Transformer 88.91 82.23 85.49 79.61 87.37 83.31 91.85 58.11 71.18 72.51 97.32 83.10 68.35 94.72 79.40 80.50
Stationary Non-stationary-transformers 78.51 87.98 82.97 86.86 68.63 76.68 90.62 55.74 69.02 89.26 95.42 92.24 98.17 96.30 97.23 83.63
DLinear DLinear 75.91 84.02 79.76 89.68 75.31 81.87 89.87 53.79 67.30 92.26 93.05 92.66 98.65 94.70 96.64 83.64
LightTS LightTS 87.10 78.42 82.53 82.40 75.78 78.95 92.58 55.27 69.21 91.98 94.72 93.33 98.37 95.97 97.15 84.23
ETSformer ETSformer 87.44 79.23 83.13 85.13 84.93 85.03 92.25 55.75 69.50 90.02 80.36 84.91 99.31 85.28 91.76 82.87
FEDformer FEDformer 72.82 81.68 76.99 90.72 75.41 82.36 90.47 58.10 70.76 99.95 65.55 79.18 99.98 81.92 90.05 79.46
PatchTST PatchTST 87.26 82.14 84.62 88.34 70.96 78.70 90.64 55.46 68.82 91.10 80.94 85.72 98.84 93.47 96.08 82.79
TimesNet TimesNet 88.07 80.97 84.37 88.83 74.68 81.14 89.98 56.02 69.05 91.99 93.24 92.61 98.46 95.70 97.06 84.85
GPT4TS GPT4TS 87.68 81.52 84.49 82.09 81.97 82.03 90.12 55.70 68.85 92.12 93.09 92.60 98.36 95.85 97.09 85.01
FedTADBench FedTADBench 87.31 80.06 83.53 77.69 69.37 84.09 90.49 57.44 70.27 90.63 84.43 87.42 97.67 94.41 96.01 84.26
PeFAD PeFAD 88.64 82.05 85.22 73.42 87.31 78.94 89.89 62.39 73.66 88.71 89.78 88.73 96.93 95.94 96.43 84.60
InterFusion InterFusion 84.06 83.52 83.78 84.83 78.49 81.54 90.66 58.11 70.82 96.76 78.45 86.65 83.61 83.45 83.53 81.26
Peri-midFormer Peri-midFormer 86.97 81.37 84.08 88.66 74.02 80.68 90.02 54.03 67.53 90.74 92.55 91.64 98.46 94.06 96.21 84.03
UniTS UniTS 82.42 84.99 83.69 91.32 73.04 81.16 90.58 62.55 74.00 92.60 92.42 92.51 98.45 96.19 97.31 85.73
SymTime (Ours) 88.08 83.37 85.66 89.46 75.31 81.77 91.06 61.51 73.43 95.94 91.39 93.61 98.90 95.36 97.10 86.31
Table 31: The full fine-tuning results of the time series long-term forecasting task under different sizes of pre-training datasets. The brief results of this experiment are shown in Table 1. Red: best, Blue: second best.
{footnotesize }
Methods 0B 1B 10B 25B 50B
[…]
96 0.335 0.376 0.324 0.355 0.319 0.353 0.321 0.358 0.318 0.353
192 0.404 0.382 0.363 0.385 0.374 0.382 0.369 0.383 0.362 0.380
336 0.419 0.443 0.392 0.410 0.391 0.412 0.394 0.407 0.386 0.402
720 0.446 0.435 0.426 0.440 0.422 0.423 0.426 0.423 0.419 0.423
ETTm1 Avg 0.401 0.409 0.376 0.398 0.376 0.393 0.378 0.393 0.371 0.390
96 0.181 0.271 0.195 0.257 0.183 0.265 0.177 0.264 0.174 0.257
192 0.241 0.323 0.268 0.316 0.249 0.310 0.245 0.302 0.238 0.299
336 0.334 0.361 0.310 0.350 0.295 0.346 0.297 0.338 0.295 0.337
720 0.417 0.401 0.394 0.400 0.398 0.396 0.391 0.396 0.390 0.392
ETTm2 Avg 0.293 0.339 0.292 0.331 0.281 0.329 0.278 0.325 0.274 0.321
96 0.422 0.428 0.402 0.411 0.386 0.404 0.384 0.402 0.376 0.400
192 0.457 0.459 0.447 0.462 0.430 0.432 0.434 0.436 0.428 0.431
336 0.523 0.494 0.489 0.467 0.473 0.463 0.468 0.457 0.463 0.456
720 0.547 0.515 0.507 0.495 0.486 0.478 0.481 0.479 0.450 0.458
ETTh1 Avg 0.487 0.474 0.461 0.459 0.444 0.444 0.434 0.438 0.430 0.436
96 0.298 0.350 0.302 0.355 0.295 0.357 0.301 0.352 0.293 0.348
192 0.373 0.403 0.364 0.416 0.369 0.405 0.369 0.403 0.364 0.397
336 0.401 0.434 0.458 0.442 0.387 0.424 0.389 0.426 0.385 0.423
720 0.433 0.461 0.490 0.464 0.454 0.447 0.427 0.438 0.420 0.441
ETTh2 Avg 0.376 0.412 0.403 0.419 0.376 0.408 0.371 0.405 0.365 0.402
96 0.185 0.221 0.173 0.219 0.170 0.218 0.175 0.217 0.166 0.213
192 0.217 0.279 0.223 0.259 0.232 0.263 0.221 0.263 0.212 0.254
336 0.276 0.304 0.285 0.294 0.266 0.296 0.271 0.297 0.267 0.294
720 0.353 0.353 0.347 0.357 0.332 0.340 0.345 0.350 0.342 0.344
Weather Avg 0.257 0.289 0.257 0.282 0.250 0.279 0.253 0.282 0.247 0.276
96 0.162 0.254 0.168 0.263 0.166 0.262 0.175 0.272 0.162 0.253
192 0.179 0.268 0.182 0.270 0.173 0.269 0.183 0.276 0.173 0.264
336 0.217 0.288 0.196 0.286 0.209 0.300 0.205 0.295 0.194 0.285
720 0.216 0.324 0.250 0.321 0.236 0.312 0.216 0.308 0.220 0.304
ECL Avg 0.193 0.284 0.199 0.285 0.196 0.286 0.195 0.288 0.187 0.276
96 0.460 0.288 0.451 0.289 0.438 0.293 0.442 0.282 0.432 0.280
192 0.452 0.287 0.461 0.291 0.450 0.292 0.452 0.289 0.444 0.287
336 0.461 0.330 0.469 0.296 0.472 0.294 0.467 0.296 0.458 0.293
720 0.510 0.336 0.510 0.336 0.508 0.314 0.502 0.309 0.492 0.303
Traffic Avg 0.471 0.310 0.473 0.303 0.473 0.294 0.467 0.299 0.457 0.291
96 0.120 0.235 0.087 0.203 0.098 0.208 0.087 0.205 0.084 0.201
192 0.188 0.298 0.192 0.309 0.186 0.301 0.177 0.297 0.174 0.295
336 0.346 0.428 0.341 0.423 0.337 0.420 0.326 0.414 0.331 0.416
720 0.878 0.699 0.861 0.706 0.850 0.701 0.838 0.689 0.847 0.694
Exchange Avg 0.383 0.415 0.370 0.410 0.368 0.407 0.357 0.401 0.359 0.401
Average 0.358 0.366 0.354 0.361 0.345 0.355 0.342 0.354 0.336 0.349
Table 32: The full fine-tuning results of the time series short-term forecasting task under different sizes of pre-training datasets. The brief results of this experiment are shown in Table 2. Red: best, Blue: second best.
{footnotesize }
Methods Metric 0B 1B 10B 25B 50B
SMAPE 13.291 13.341 13.332 13.380 13.355
MASE 2.981 2.986 2.985 3.012 2.997
Yearly OWA 0.782 0.784 0.783 0.788 0.786
SMAPE 10.270 10.274 10.197 10.228 10.060
MASE 1.224 1.218 1.212 1.219 1.183
Quartly OWA 0.913 0.911 0.905 0.909 0.872
SMAPE 13.545 12.811 12.833 12.662 12.608
MASE 1.053 0.955 0.959 0.932 0.925
Monthly OWA 0.964 0.893 0.896 0.877 0.872
SMAPE 5.186 5.070 5.003 5.034 4.941
MASE 3.498 3.479 3.350 3.372 3.327
Others OWA 1.097 1.082 1.055 1.061 1.045
SMAPE 12.283 11.937 11.924 11.862 11.785
MASE 1.660 1.611 1.605 1.601 1.584
Avg. OWA 0.887 0.861 0.859 0.856 0.849
Table 33: The full fine-tuning results of the time series classification task under different sizes of pre-training datasets. The brief results of this experiment are shown in Figure 7 (a). Red: best, Blue: second best.
{footnotesize }
Datasets 0B 1B 10B 25B 50B
EthanolConcentration 33.08 30.04 34.22 35.74 37.30
FaceDetection 51.31 58.12 58.12 58.12 69.20
Handwriting 35.76 36.24 36.24 36.00 36.70
Heartbeat 70.24 71.71 72.20 73.17 74.15
JapaneseVowels 94.86 95.95 94.86 96.76 98.11
PEMS-SF 86.71 88.44 88.44 92.49 97.11
SelfRegulationSCP1 86.69 85.67 87.71 88.05 89.76
SelfRegulationSCP2 56.11 57.78 57.78 58.89 58.89
SpokenArabicDigits 89.77 93.91 95.91 97.04 98.86
UWaveGestureLibrary 78.75 85.63 85.63 87.19 89.38
Average Accuracy 68.33 70.35 71.11 71.34 74.90
Table 34: The full fine-tuning results of the time series imputation task under different sizes of pre-training datasets. The brief results of this experiment are shown in Table 2. Red: best, Blue: second best.
{footnotesize }
Methods 0B 1B 10B 25B 50B
Mask Ratio MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
12.5% 0.034 0.111 0.034 0.112 0.032 0.111 0.033 0.110 0.032 0.110
25% 0.050 0.125 0.042 0.116 0.041 0.118 0.036 0.116 0.034 0.113
37.5% 0.040 0.122 0.037 0.119 0.039 0.118 0.036 0.118 0.037 0.118
50% 0.044 0.128 0.042 0.129 0.042 0.128 0.042 0.126 0.041 0.126
ETTm1 Avg 0.042 0.122 0.039 0.119 0.038 0.119 0.037 0.118 0.036 0.117
12.5% 0.030 0.088 0.027 0.087 0.028 0.087 0.026 0.086 0.024 0.084
25% 0.038 0.101 0.029 0.087 0.029 0.087 0.028 0.088 0.024 0.086
37.5% 0.040 0.120 0.032 0.107 0.028 0.096 0.027 0.091 0.027 0.089
50% 0.045 0.114 0.036 0.109 0.034 0.110 0.032 0.106 0.030 0.093
ETTm2 Avg 0.038 0.106 0.031 0.097 0.030 0.095 0.028 0.093 0.026 0.088
12.5% 0.100 0.208 0.085 0.185 0.087 0.186 0.080 0.180 0.074 0.179
25% 0.120 0.238 0.116 0.222 0.106 0.221 0.103 0.207 0.082 0.190
37.5% 0.118 0.234 0.118 0.227 0.100 0.206 0.103 0.209 0.100 0.205
50% 0.140 0.241 0.133 0.235 0.137 0.240 0.130 0.232 0.123 0.230
ETTh1 Avg 0.112 0.230 0.113 0.217 0.107 0.213 0.104 0.207 0.095 0.201
12.5% 0.061 0.158 0.061 0.158 0.057 0.151 0.054 0.146 0.051 0.138
25% 0.066 0.161 0.058 0.153 0.059 0.155 0.058 0.154 0.055 0.146
37.5% 0.065 0.158 0.070 0.164 0.067 0.160 0.059 0.155 0.059 0.152
50% 0.068 0.164 0.073 0.167 0.068 0.166 0.066 0.161 0.064 0.157
ETTh2 Avg 0.065 0.160 0.066 0.160 0.063 0.158 0.059 0.154 0.057 0.148
12.5% 0.046 0.123 0.039 0.123 0.039 0.123 0.038 0.123 0.037 0.122
25% 0.046 0.148 0.049 0.140 0.048 0.139 0.047 0.139 0.046 0.139
37.5% 0.062 0.168 0.063 0.163 0.062 0.161 0.060 0.161 0.060 0.160
50% 0.076 0.181 0.076 0.182 0.075 0.182 0.076 0.183 0.075 0.181
ECL Avg 0.058 0.155 0.057 0.152 0.056 0.151 0.055 0.152 0.055 0.151
12.5% 0.029 0.044 0.030 0.046 0.025 0.036 0.026 0.038 0.025 0.035
25% 0.038 0.054 0.030 0.044 0.036 0.052 0.029 0.042 0.027 0.037
37.5% 0.038 0.058 0.037 0.057 0.034 0.052 0.032 0.045 0.029 0.039
50% 0.040 0.057 0.036 0.054 0.036 0.052 0.034 0.046 0.032 0.042
Weather Avg 0.036 0.053 0.033 0.050 0.033 0.048 0.030 0.043 0.028 0.038
Average 0.058 0.138 0.057 0.132 0.055 0.131 0.052 0.128 0.049 0.124
Table 35: The full fine-tuning results of the time series classification task under different sizes of pre-training datasets. The brief results of this experiment are shown in Figure 7 (b). Red: best, Blue: second best.
{footnotesize }
Datasets 0B 1B 10B 25B 50B
Metrics P R F1 P R F1 P R F1 P R F1 P R F1
SMD 86.70 80.72 83.60 88.08 83.37 85.66 87.46 81.05 84.13 86.89 82.02 84.39 88.08 83.37 85.66
MSL 89.28 73.68 80.74 89.13 73.59 80.62 89.48 74.59 81.36 89.34 73.91 80.90 89.46 75.31 81.77
SMAP 89.97 54.16 67.61 90.06 53.56 67.17 90.08 54.51 67.92 90.11 54.63 68.02 91.06 61.51 73.43
SWaT 91.57 85.45 88.40 92.21 92.70 92.45 92.23 92.77 92.50 92.28 92.87 92.57 95.94 91.39 93.61
PSM 98.69 94.24 96.41 98.53 94.30 96.37 98.65 94.41 96.48 98.79 94.13 96.40 91.39 93.61 98.90
Avg. 91.24 77.65 83.40 91.60 79.50 84.45 91.58 79.47 84.48 91.48 79.51 84.46 95.36 97.10 86.31
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We made our main claims in the abstract and introduction. We have clearly pointed out the problem to be solved and the method to be used in the abstract and introduction.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The $S^{2}$ bimodal data generation mechanism is the core method and contribution of this paper. In Appendix B.9, we further explore the current limitations in this aspect and will try to improve it in subsequent work.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced. The core ideas of this paper are further explained and proved by Takens’s theorem and symbolic dynamics in the beginning Section 3.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We describe the proposed methodology in detail in Section 3 and provide further explanation in Appendix B. Experimental details are outlined in Appendix C. Additionally, all the code related to the proposed method is included in the supplementary material.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The code is available in the supplementary material.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We describe the datasets used in detail in Appendix C.1, and further elaborate on the model pre-training and fine-tuning configurations for downstream tasks in Appendix C.3 and C.4.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: In Section 4.4 we conducted multiple experiments in the ablation experiments and drew error bars in the bar graphs to further ensure the reliability of the ablation conclusions. However, due to the large number of experiments, detailed standard deviations are not shown. Instead, we uniformly present them in the headings of each table in Appendix F.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: In Appendix C we describe the hardware environment in detail. Using the pre-trained configuration in Appendix C.3 on our eight A6000s is enough to fill the memory of all devices. The computing resources and memory required for fine-tuning the model are shown in Figure 5 (right).
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We have read the NeurIPS Code of Ethics and conducted it in the paper conform in every respect.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We illustrate broader impacts of our research in Appendix G.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The models compared and the datasets used in this article are open source, and the corresponding papers or web pages are cited.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We provide the code for the proposed methodology in the supplementary material and will make it publicly available at an appropriate time.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The LLM does not affect the core methodology, scientific rigor or originality of the research.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v2

@@
 B.4 Analysis of Existing Large-scale Datasets for Time Series Pre-training
+(a) data scarcity
+(b) data imbalance
+Figure 13: The scarcity and imbalance of time series pre-training dataset (taking the largest open-source time series dataset Time-300BB as an example Time-MoE ).
+(a) time series datasets are data-scarce compared to text datasets in natural language processing and video understanding datasets in computer vision.
+(b) Large-scale time series pre-training datasets face serious distribution imbalance problems.
 Table 11: Time-300B time series dataset from Time-MoE Time-MoE .
@@
 This paper surveys the pre-training datasets used by the three current mainstream pre-trained foundation models—Time-MoE Time-MoE , Moirai MOIRAI , and Timer Timer —as well as the datasets utilized in the study of time series scaling laws Time-Scaling-Laws , which are shown in Tables 11, 12, 13 and 14.
+In Figure 13 (a) we demonstrate that the current largest time series datasets are still smaller than those in CV and NLP.
 Imbalanced domain distribution issues in large-scale time series datasets.
 The distribution of data across various domains indicates that the four large-scale time series pre-training datasets all face issues with imbalanced domain data distribution.
-For instance, domains such as Nature, Energy and Transport have the most datasets Informer , while others like Sales, IoT, Web, Finance and Multiple Sources suffer from extremely low data volumes due to difficulties in data collection or data privacy concerns.
+For instance, domains such as Nature, Energy and Transport have the most datasets Informer , while others like Sales, IoT, Web, Finance and Multiple Sources suffer from extremely low data volumes due to difficulties in data collection or data privacy concerns as shown in Figure 13 (b).
 According to the scaling laws of neural networks, the imbalance in the pre-training dataset distribution can lead to significant performance biases in in-domain and out-of-domain forecasting tasks for the trained foundation models Time-Scaling-Laws ; graph-forecasting , meaning there is a considerable performance gap between domains with less data and those with more data.
@@
 As SymTime adds masks in units of patches of length 16 during pre-training, we also add masks in the form of 16-length patches.
-The reconstruction effect of the masked parts by the time series encoder is shown in Figure 13 and 14.
+The reconstruction effect of the masked parts by the time series encoder is shown in Figure 14 and 15.
 $S^{2}$ Dataset Out-of-Domain Data.
@@
 Avg. 91.24 77.65 83.40 91.60 79.50 84.45 91.58 79.47 84.48 91.48 79.51 84.46 95.36 97.10 86.31
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4. Justification: We made our main claims in the abstract and introduction. We have clearly pointed out the problem to be solved and the method to be used in the abstract and introduction.
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes]
-  9.
-Justification: The $S^{2}$ bimodal data generation mechanism is the core method and contribution of this paper.
-In Appendix B.9, we further explore the current limitations in this aspect and will try to improve it in subsequent work.
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory assumptions and proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [Yes]
-  14.
-Justification: All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-The core ideas of this paper are further explained and proved by Takens’s theorem and symbolic dynamics in the beginning Section 3.
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental result reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes]
-  19.
-Justification: We describe the proposed methodology in detail in Section 3 and provide further explanation in Appendix B.
-Experimental details are outlined in Appendix C.
-Additionally, all the code related to the proposed method is included in the supplementary material.
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [Yes]
-  24. Justification: The code is available in the supplementary material.
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental setting/details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes]
-  29.
-Justification: We describe the datasets used in detail in Appendix C.1, and further elaborate on the model pre-training and fine-tuning configurations for downstream tasks in Appendix C.3 and C.4.
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment statistical significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [Yes]
-  34.
-Justification: In Section 4.4 we conducted multiple experiments in the ablation experiments and drew error bars in the bar graphs to further ensure the reliability of the ablation conclusions.
-However, due to the large number of experiments, detailed standard deviations are not shown.
-Instead, we uniformly present them in the headings of each table in Appendix F.
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments compute resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [Yes]
-  39.
-Justification: In Appendix C we describe the hardware environment in detail.
-Using the pre-trained configuration in Appendix C.3 on our eight A6000s is enough to fill the memory of all devices.
-The computing resources and memory required for fine-tuning the model are shown in Figure 5 (right).
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code of ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes]
-  44. Justification: We have read the NeurIPS Code of Ethics and conducted it in the paper conform in every respect.
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [Yes]
-  49. Justification: We illustrate broader impacts of our research in Appendix G.
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [N/A]
-  54. Justification: The paper poses no such risks.
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes]
-  59. Justification: The models compared and the datasets used in this article are open source, and the corresponding papers or web pages are cited.
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [Yes]
-  64. Justification: We provide the code for the proposed methodology in the supplementary material and will make it publicly available at an appropriate time.
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and research with human subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [N/A]
-  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional review board (IRB) approvals or equivalent for research with human subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [N/A]
-  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
-  76. 16.
-    Declaration of LLM usage
-  77.
-Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
-Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
-  78. Answer: [N/A]
-  79. Justification: The LLM does not affect the core methodology, scientific rigor or originality of the research.
-  80. Guidelines:
-    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
-    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v3

@@
 2 Related Work
-Pre-trained foundation models (PTFMs) Time-LLM ; CLIP ; BERT ; BEiT have been demonstrated to adapt to a variety of downstream tasks after fine-tuning on specific datasets, exhibiting excellent generalization and scalability Chronos ; TimeMIL .
+Pre-trained foundation models (PTFMs) Time-LLM ; CLIP ; BERT ; BEiT ; ijcai2025p1067 have been demonstrated to adapt to a variety of downstream tasks after fine-tuning on specific datasets, exhibiting excellent generalization and scalability Chronos ; TimeMIL .
 Inspired by this, recent years have seen significant progress in PTFMs for TSA Transformer-in-TSA ; KDD-Survey , with the emergence of various pre-training methods.
@@
 Series–Symbol Contrastive Learning.
-To leverage series-symbol data pairs with strong correlations, we employ contrastive learning to enhance time series representation using symbolic information.
+To leverage series-symbol data pairs with strong correlations (The correspondence between positive and negative samples is shown in Appendix C.2), we employ contrastive learning to enhance time series representation using symbolic information.
 Using momentum encoders MoCo_v1 , we project [CLS] embeddings via linear projections $g_{t},g_{s}$ and define $\mathrm{sim}(t,s)=g_{t}(t_{\mathrm{cls}})^{\top}g^{\prime}_{s}(s^{\prime}_{\mathrm{cls}})$, where $g^{\prime}_{s}(s^{\prime}_{\text{cls}})$ is the normalized symbol features generated by the momentum model.
@@
   • Classification: patch the series, encode, and classify via a linear head PatchTST ; TimeXer .
-  • Reconstruction (forecasting, imputation, anomaly detection) Peri-midFormer : decompose each series into trend and periodic components; regress trend directly, patch and encode the periodic part, then recombine for the final output.
+  • Reconstruction (forecasting, imputation, anomaly detection) Peri-midFormer ; Sub-Adjacent : decompose each series into trend and periodic components; regress trend directly, patch and encode the periodic part, then recombine for the final output.
 4 Experiments
@@
 We use mean squared error (MSE) and mean absolute error (MAE) as the metrics for long-term forecasting and imputation tasks; overall weighted average (OWA) for short-term forecasting, which is unique metrics for M4 benchmark M4team2018dataset ; accuracy for classification; precision, recall and F1 score for anomaly detection.
-Detailed descriptions of datasets and metrics for each task are provided in Appendix C.1 and C.2, while the pre-training and fine-tuning configurations for SymTime across downstream tasks are outlined in Appendix C.3 and C.4.
+Detailed descriptions of datasets and metrics for each task are provided in Appendix C.1 and C.3, while the pre-training and fine-tuning configurations for SymTime across downstream tasks are outlined in Appendix C.4 and C.5.
 Baselines.
@@
 Acknowledgement
-This work was supported in part by Natural Science Basic Research Program of Shaanxi under Grant 2025JC-QYCX-060, in part by the National Natural Science Foundation of China under Grant 62206205, 62201419 and 62471371, in part by the Young Talent Fund of Association for Science and Technology in Shaanxi, China under Grant 20230129, in part by the Guangdong High-level Innovation Research Institution Project under Grant 2021B0909050008, in part by the Guangzhou Key Research and Development Program under Grant 202206030003, and in part by the Fundamental Research Funds for the Central Universities under Grant QTZX25111.
+This work was supported in part by Natural Science Basic Research Program of Shaanxi under Grant 2025JC-QYCX-060, in part by the National Natural Science Foundation of China under Grant 62206205, 62201419, 62472345, and 62471371, in part by the Young Talent Fund of Association for Science and Technology in Shaanxi, China under Grant 20230129, in part by the Guangdong High-level Innovation Research Institution Project under Grant 2021B0909050008, in part by the Guangzhou Key Research and Development Program under Grant 202206030003, and in part by the Fundamental Research Funds for the Central Universities under Grant QTZX25111.
 This research was supported by Huawei’s Al Hundred Schools Program and was carried out using the Huawei Ascend AI technology stack.
@@
 For anomaly detection in time series, we experiment with five datasets: SMD SMD , MSL MSL , SMAP MSL , SWaT SWaT , and PSM PSM .
-C.2 Metrics
+C.2 Correspondence between Positive and Negative Samples in SymTime
+In our $S^{2}$ data generation mechanism, we randomly construct a symbolic expression $f(\cdot)$ and an input stimulus time series $X$, and then forward propagate the symbolic expression to obtain the complex system’s response time series $Y$.
+Therefore, the generated time series $Y$ has a natural generating-being relationship with the symbolic expression $f(\cdot)$.
+Since symbolic expressions can model complex systems, and different symbolic expressions can correspond to different symbolic systems, the generated time series can also reflect the corresponding complex system’s representation.
+When performing contrastive learning, we consider the time series and the symbolic expression that generated it as positive samples, and other unrelated symbols as negative samples.
+C.3 Metrics
 We assess the five TSA tasks using various metrics.
@@
 $\mathbf{X}_{i}$ means the $i$-th future time point.
-C.3 Pre-training
+C.4 Pre-training
 Model Architecture.
@@
 Encoder Layers $d_{\mathrm{model}}$ $d_{\mathrm{ff}}$ Heads Params
-Time 6 512 2048 8 19M
-Symbol 6 786 3072 12 67M
+Time Series 6 512 2048 8 19M
+Symbol 6 768 3072 12 67M
 Model Hyper-parameter.
@@
 Due to the large size of our generat $S^{2}$ dataset, we load data into the GPU in batches during each epoch for pre-training.
-C.4 Fine-tuning
+C.5 Fine-tuning
 For the five major tasks in TSA, we conduct downstream task fine-tuning experiments using the configurations in Table 17.
@@
 Anomaly Detection 3, 6 $10^{-4}-5\times 10^{-4}$ MSE 4-64 12
-C.5 Ablation Experiments Details
+C.6 Ablation Experiments Details
 Ablation study on pre-training strategies and objectives.
@@
     w/o Distill: The contrastive loss in pre-training does not use the pseudo objective of momentum distillation.
-C.6 Ablation Experiments on Short-term Forecasting
+C.7 Ablation Experiments on Short-term Forecasting
 Setup.
@@
 Figure 17 (a) shows that as the size of the pre-training dataset increases, SymTime’s performance on the Quarterly dataset improves significantly.
-C.7 The Ablation of Backbone in SymTime
+C.8 The Ablation of Backbone in SymTime
 Setup.
@@
 Therefore, if the basic pre-training requirements are met (pre-training loss can be successfully optimized), a more lightweight model can be used for fine-tuning on downstream tasks.
-C.8 The Impact and Ablation of Pre-interpolation on Time Series Imputation Task
+C.9 The Impact and Ablation of Pre-interpolation on Time Series Imputation Task
 SymTime adds masks randomly at the patch level during pre-training for time series reconstruction.
@@
 In CV and NLP CLIP , PTFMs have been demonstrated to adapt to a variety of downstream tasks after fine-tuning on specific datasets, exhibiting excellent generalization and scalability.
-Inspired by this, recent years have seen significant progress in PTFMs for TSA Transformer-in-TSA ; KDD-Survey ; TimeDiT , with the emergence of various pre-training methods.
+Inspired by this, recent years have seen significant progress in PTFMs for TSA Transformer-in-TSA ; KDD-Survey ; TimeDiT ; FreEformer , with the emergence of various pre-training methods.
 MOIRAI, through MTM and reconstruction, has been pre-trained on large datasets (27B), yielding a universal forecasting model with significant zero-shot advantages MOIRAI .
@@
 Avg. 91.24 77.65 83.40 91.60 79.50 84.45 91.58 79.47 84.48 91.48 79.51 84.46 95.36 97.10 86.31
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: We made our main claims in the abstract and introduction. We have clearly pointed out the problem to be solved and the method to be used in the abstract and introduction.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9.
+Justification: The $S^{2}$ bimodal data generation mechanism is the core method and contribution of this paper.
+In Appendix B.9, we further explore the current limitations in this aspect and will try to improve it in subsequent work.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [Yes]
+  14.
+Justification: All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+The core ideas of this paper are further explained and proved by Takens’s theorem and symbolic dynamics in the beginning Section 3.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19.
+Justification: We describe the proposed methodology in detail in Section 3 and provide further explanation in Appendix B.
+Experimental details are outlined in Appendix C.
+Additionally, all the code related to the proposed method is included in the supplementary material.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Justification: The code is available in the supplementary material.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29.
+Justification: We describe the datasets used in detail in Appendix C.1, and further elaborate on the model pre-training and fine-tuning configurations for downstream tasks in Appendix C.4 and C.5.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+  34.
+Justification: In Section 4.4 we conducted multiple experiments in the ablation experiments and drew error bars in the bar graphs to further ensure the reliability of the ablation conclusions.
+However, due to the large number of experiments, detailed standard deviations are not shown.
+Instead, we uniformly present them in the headings of each table in Appendix F.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39.
+Justification: In Appendix C we describe the hardware environment in detail.
+Using the pre-trained configuration in Appendix C.4 on our eight A6000s is enough to fill the memory of all devices.
+The computing resources and memory required for fine-tuning the model are shown in Figure 5 (right).
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: We have read the NeurIPS Code of Ethics and conducted it in the paper conform in every respect.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [Yes]
+  49. Justification: We illustrate broader impacts of our research in Appendix G.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: The paper poses no such risks.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: The models compared and the datasets used in this article are open source, and the corresponding papers or web pages are cited.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [Yes]
+  64. Justification: We provide the code for the proposed methodology in the supplementary material and will make it publicly available at an appropriate time.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  78. Answer: [N/A]
+  79. Justification: The LLM does not affect the core methodology, scientific rigor or originality of the research.
+  80. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.